### v2

# When Models Outthink Their Safety: Mitigating Self-Jailbreak in Large Reasoning Models with Chain-of-Guardrails

Abstract
Large Reasoning Models (LRMs) demonstrate remarkable capabilities on complex reasoning tasks but remain vulnerable to severe safety risks, including harmful content generation and jailbreak attacks. Existing mitigation strategies rely on injecting heuristic safety signals during training, which often suppress reasoning ability and fail to resolve the safety-reasoning trade-off. To systematically investigate this issue, we analyze the reasoning trajectories of diverse LRMs and uncover a phenomenon we term Self-Jailbreak, where models override their own risk assessments and justify responding to unsafe prompts. This finding reveals that LRMs inherently possess the ability to reject unsafe queries, but this ability is compromised, resulting in harmful outputs. Building on these insights, we propose the Chain-of-Guardrail (CoG), a training framework that recomposes or backtracks unsafe reasoning steps, steering the model back onto safe trajectories while preserving valid reasoning chains. Extensive experiments across multiple reasoning and safety benchmarks demonstrate that CoG substantially improves the safety of current LRMs while preserving comparable reasoning ability, significantly outperforming prior methods that suffer from severe safety–reasoning trade-offs.
1 Introduction
Large Reasoning Models (LRMs), such as OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Qwen3 (Yang et al., 2025), have demonstrated remarkable capabilities in solving complex tasks and are increasingly being adopted across a wide range of domains. Despite these advances, researchers have identified that current LRMs remain hindered by significant safety vulnerabilities, such as the generation of harmful content as well as the susceptibility to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025; Wang et al., 2025a). These issues highlight a critical obstacle to the reliable and responsible use of LRMs (Weidinger et al., 2021). Therefore, addressing these challenges is critical for ensuring that LRMs operate safely and remain aligned with human values.
While recent efforts (Wang et al., 2025b; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have attempted to mitigate these risks, they largely exhibit a pronounced safety-reasoning trade-off. Specifically, current studies have primarily focused on curating safety-oriented datasets based on heuristic rules and fine-tuning LRMs to produce refusal responses. Unfortunately, while such approaches can improve safety to some extent, the pattern-based behavior often conflicts with the flexible and multi-step deliberation required for complex reasoning, thereby substantially degrading reasoning ability. Therefore, it is critical to identify the underlying causes of the unsafe behavior in LRMs, and develop more principled strategies that can enhance safety without sacrificing reasoning performance.
To this end, we conduct a systematic analysis of LRMs’ reasoning trajectories and identify a critical failure phenomenon, which we term Self-Jailbreak. Specifically, as demonstrated in Figure 1, we decompose LRM’s reasoning thinking trajectory into three stages: Risk Awareness, Risk Analysis, and Response Strategy. We observe that LRMs often detect harmful intent during the Risk Awareness stage, but subsequently override this recognition during the Risk Analysis stage, and ultimately persuade themselves to respond to unsafe queries. This reveals that LRMs possess an inherent ability to reject unsafe queries, yet it is undermined by Self-Jailbreak, resulting in unsafe generation. Furthermore, we categorize Self-Jailbreak into four distinct types—Benign Reframing, Logical Fallacy, Warning, and Risk Misrecognition—and find that Warning accounts for most unsafe outputs.
Existing approaches that solely inject heuristic safety signals during training fail to resolve the safety–reasoning trade-off. While they suppress overtly unsafe outputs, such interventions do not address unsafe reasoning steps embedded within reasoning trajectories. Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving reasoning quality. Specifically, CoG leverages precise classification of Self-Jailbreak to perform targeted safety interventions, mitigating unsafe behavior while preserving the model’s reasoning abilities and reasoning patterns. To achieve this, CoG employs two strategies: 1) Safety Recomposition, which systematically reorganizes unsafe reasoning into entirely new, safety-assured chains, and 2) Safety Backtrack, which preserves the model’s original reasoning path while incorporating self-check and backtracking mechanisms to correct risky steps before they lead to unsafe conclusions. The resulting safety-oriented reasoning chains are then used to fine-tune LRMs, aligning them toward safe responses while maintaining the reasoning capability.
Figure 1: Illustration of “Self-Jailbreak.” The Instruct model (left) directly refuses harmful queries, whereas the LRM (right) identifies risk during Risk Awareness but later engages in Self-Jailbreak (red text) during Risk Analysis, ultimately overriding its initial safety judgment and producing harmful content in the Response Strategy stage.
Extensive experiments across multiple safety and reasoning benchmark datasets demonstrate that COG excels at balancing the safety and reasoning performance of LRMs. On Qwen3-32B, CoG improves the reasoning ability by 13.5% over SafeKey while achieving state-of-the-art performance on GPQA and AIME (Rein et al., 2024; Mathematical Association of America (2024), MAA), maintaining comparable reasoning performance. Further analysis of reasoning trajectories reveals that, unlike other approaches, CoG preserves the model’s intrinsic reasoning patterns, confirming that its safety gains do not come at the expense of the model’s original deliberation trajectory. In summary, our contributions include: In summary, our contributions include:
  • To the best of our knowledge, we are the first to identify and characterize the Self-Jailbreak phenomenon, revealing it as a key factor behind the poor safety performance of LRMs.
  • We propose a systematic analytical framework to categorize Self-Jailbreak instances and quantify their impact on model behavior, finding that Benign Framing and Warning constitute the majority of the Self-Jailbreak phenomenon.
  • We introduce the Chain-of-Guardrail (CoG), a training framework that mitigates Self-Jailbreak and substantially improves model safety while preserving reasoning capabilities. Experiment results show that CoG improves the safety-reasoning trade-off by 3.39%, reaching the state-of-the-art (SOTA) across multiple LRMs.
Figure 2: Processing of harmful queries by LRM, illustrating the workflow from receiving prompts, through risk detection during reasoning, to the four possible outcomes: aware of risk and responds, aware of risk and refuses, unaware of risk and responds, unaware of risk and refuses.
2 Motivation: Preliminary Experiments On Self-Jailbreak
This section examines the thinking trajectory of LRMs, drawing on pilot experimental results (reported in Appendix B.1) which indicate that most LRMs become unsafe when the thinking mode is enabled. Subsequently, we find that even if LRMs are capable of recognizing latent dangerous intentions, they still respond to the harmful query. Such results shed light on our hypothesis: Self-Jailbreak. We further validate the existence of Self-Jailbreak with experiments, which represent two significant findings: (1) Most LRMs encounter Self-Jailbreak during the reasoning trajectory. (2) Among the different categories of Self-Jailbreak, Warning is the primary source.
2.1 Hypothesis: Self-Jailbreak via Thinking–Answer Misalignment
We analyzed the models’ chains of thought, discovering that their thinking trajectory can be decomposed into three stages: risk awareness, risk analysis, and response strategy, which enables a deeper examination of how thinking affects the safety of LRMs.
Building on the previous experiment, we employ LLM-as-Judge (Team, 2024; Gu et al., 2024) to assess whether the LRM demonstrates awareness of potential risks during the risk awareness stage, and use Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content. This setup enables us to examine potential misalignments between the model’s reasoning trajectory and its generated answers. As illustrated in Figure 2, our analysis revealed a striking contradiction:
LRMs are able to identify most of the harmful queries but still generate unsafe answers. 94.2% of harmful queries, as well as users’ potentially malicious intents, are recognized; however, 52.30% of their generated outputs still contain hazardous content. In contrast, only 39.90% of the cases reflect awareness of risk followed by a safe response. This inconsistency between “thinking” and “response” raises a critical concern: during subsequent reasoning, LRMs may tend to persuade themselves to respond to such queries. Based on this observation, we propose the following hypothesis:
Self-Jailbreak During the thinking process, LRMs often exhibit a phenomenon we term Self-Jailbreak, where models persuade themselves to produce responses to potentially harmful requests in an attempt to assist the user, even after recognizing the risky nature of the queries and the user’s underlying malicious intent.
Table 1: Self-Jailbreak category distribution of various LRMs (%)
Model Response Safety Self-Jailbreak Categories
Safe Unsafe Benign Reframing Warning Logical Fallacies Harm Identification
DS-R1-0528 4.07 95.93 42.29 52.12 0.51 1.02
DS-Llama-70B 0.50 99.50 47.89 45.67 0.91 5.04
Qwen3-8B 0.21 99.79 38.60 50.69 2.41 8.10
Qwen3-14B 0.15 99.85 37.30 53.10 1.73 7.32
Qwen3-32B 0.36 99.64 36.32 52.51 2.20 8.61
2.2 Categorization and Existence of Self-Jailbreak
Building on this hypothesis and inspiration from backtracking during mathematical reasoning  (Chen et al., 2025), we designed a more rigorous classification framework, which is guided by the principle that the model’s reasoning trajectory should not generate content that directly or indirectly facilitates harmful user behaviors. We categorize these behaviors into four types:
Benign Reframing:
The model recognizes the risk but attempts to reinterpret the user’s intent benignly, offering help from such interpretation.
Warning:
The model assumes that appending a warning to harmful content is sufficient to prevent its misuse, and proceeds to generate its answer.
Logical Fallacies:
Given a query containing logical contradictions, the model detects potential harm but falls into logical traps, leading to logical fallacies.
Harm Identification:
The model fails to recognize the harm in the query, resulting in answering.
To investigate the prevalence of Self-Jailbreak in LRMs, we categorize and analyze the thinking trajectory of several contemporary LRMs on 2k queries in WildJailbreak (Jiang et al., 2024a). As shown in Table 1, we observe that:
(1) Self-Jailbreak is a prevalent phenomenon observed in various LRMs. Experiment results indicate that Self-Jailbreak occurs frequently across LRMs, including peripheral models like DeepSeek-R1-0528 (DeepSeek-AI, 2025), highlighting the importance and relevance of our study.
(2) Warning is the most common source of Self-Jailbreak. As illustrated in Table 1, Warning is the most categorized Self-Jailbreak type across various LRMs. This tendency likely stems from the fact that LRMs are largely trained on math and code data, which predisposes them to respond in ways that meet user expectations. Consequently, even when confronted with harmful queries, LRMs are likely to be helpful, inadvertently bypassing their own safety mechanisms by providing content alongside a safety disclaimer. Such behavior ultimately gives rise to Self-Jailbreak.
3 Chain-of-Guardrail: A Training Framework to Mitigate Self-Jailbreak
To systematically address the risks of Self-Jailbreak in LRMs, we propose a unified safe training framework: Chain-of-Guardrail (COG). As illustrated in Figure 3, the framework is designed to create an interpretable and structured pipeline that automatically removes or rejects unsafe content during the reasoning trajectory $c$, producing a safe and logically consistent response $y_{\text{safe}}$. To achieve this, COG employs two complementary methods: Safety Recomposition (SafR) and Safety Backtrack (SafB). SafR focuses on proactively restructuring and correcting the reasoning chain to eliminate potential safety risks, while SafB focuses on the reasoning trajectory when unsafe content is detected, ensuring that the model corrects errors and generates a safe response.
[…]
3.1 Optimization Goal of LRMs in COG
Specifically, we formulate the optimization objective as follows:
$f:(\pi_{0},x)\rightarrow y_{\text{safe}},\quad\text{s.t.}\quad\text{Preserve}(\text{Reasoning}_{\pi_{0}})$ (1)
where $\pi_{0}$ is the original model, $y_{\text{safe}}$ is the safe response, and $\text{Preserve}(\cdot)$ ensures maximal retention of the original reasoning.
To achieve this, we explicitly model the Chain-of-Thought (CoT) as a structured latent variable:
$c=\pi_{0}(x)=[d(x),a(x),p(x)],$ (2)
where $d(x)$, $a(x)$, and $p(x)$ represent the risk awareness, risk analysis, and response strategy.
Depending on the adopted process, the safety-oriented reasoning chain is constructed as follows:
Chain of Safety Recomposition (SafR):
$c_{\text{safe}}=\text{Merge}(d(x),\hat{a}(x),\hat{p}(x)),$ (3)
where $\hat{a}(x)$ and $\hat{p}(x)$ are the risk analysis and response strategy recomposed by the Safety Recomposition procedure.
Chain of Safety Backtrack (SafB):
$c^{\prime}_{\text{safe}}=[c,s(x)],$ (4)
where $s(x)$ denotes the model-generated self-check module appended to the original reasoning chain in the Safety Backtrack procedure.
The final safe response $y_{\text{safe}}$ is produced based on the safety-oriented reasoning chain derived from either SafR or SafB.
3.2 Selective Loss Masking Strategy
To avoid undesired shifts in the model’s predicted distribution due to safety interventions, we employ a selective loss masking strategy. Let the output sequence be $o=[c_{\text{safe}},y_{\text{safe}}]$ for SafR, and $o=[c^{\prime}_{\text{safe}},y_{\text{safe}}]$ for SafB, where $c_{\text{safe}}$ and $c^{\prime}_{\text{safe}}$ are defined as in Eq. 3 and Eq. 4, respectively.
We propose the utilization of a binary mask vector $m\in\{0,1\}^{|o|}$, to supervise distinct segments of the output sequence $o$ selectively. The primary training goal can be summarized as follows:
$\mathcal{L}_{SafR/SafB}=-\sum_{i=1}^{|o|}m_{i}\cdot\log P_{\theta}(o_{i}|o_{<i},x),$ (5)
This selective loss masking strategy allows for flexibility depending on the applied process:
Safety Recomposition: $m$ covers all tokens, optimizing both the revised reasoning chain and the safe response, allowing the model to learn the complete safe reasoning trajectory.
Safety Backtrack: $m$ covers only the appended self-check and final response, while the original chain is masked as context, preserving the model’s reasoning distribution and enhancing self-reflection and safety.
3.3 Workflow of COG
Both SafR and SafB aim to enhance safety by mitigating certain aspects of Self-Jailbreak, following a three-phase structure. While their core workflows overlap, they differ in how the reasoning chain is modified and integrated.
Phase 1: Input & Analysis (shared)
Initial Inference. The base model $\pi_{0}$ infers on the harmful prompt $x$, producing the initial reasoning chain $c=[a(x),p(x)]$ and response $y$.
Information Extraction. A structured prompt (see Appendix A.2) together with Qwen2.5-72B-Instruct (Team, 2024) extracts the risk analysis and response strategy.
Failure Classification. Given $x$, $a(x)$, and $p(x)$, Qwen2.5-72B-Instruct determines whether the reasoning violates safety principles and assigns a Self-Jailbreak type.
Phase 2: Data Construction (divergent)
SafR – Reasoning Chain Recomposition and Merging. For each identified failure type, a tailored prompt guides $\pi_{0}$ to perform reasoning chain recomposition, adjusting the risk analysis and response strategy with safety-oriented considerations to generate corrected sub-chains. Subsequently, these recomposed analyses $a(x)$ and predictions $p(x)$ are merged into a coherent, Safety-oriented Chain-of-Thought (S-COT).
SafB – Safety Backtrack Augmentation. Rather than recomposing, $\pi_{0}$ performs a reflective process on the original reasoning chain to generate a supplementary safety check $s(x)$. Carefully engineered transitional expressions ensure a natural integration between the reasoning chain and the check, resulting in a coherent S-COT.
Phase 3: Inference & Alignment (shared, with minor variation)
Safe Response Generation. Both SafR and SafB use the S-COT as the reasoning trajectory to guide $\pi_{0}$ in generating a safe response $y_{\text{safe}}$.
Alignment Training. In both methods, $\pi_{0}$ is further trained on the safe response data, yielding the aligned and safety-enhanced model $\pi_{1}$ with strong reasoning capabilities.
4 Experiment
4.1 Experiment Setting
Baselines We include four representative methods: STAR-1  (Wang et al., 2025b) and SafeChain  (Jiang et al., 2025) improve safety by performing SFT on a curated set of safe data. SAFEPATH  (Jeung et al., 2025) guides model behavior by training it to generate a fixed prefix, and SafeKey  (Zhou et al., 2025b) enforces stricter safety alignment by introducing two additional safety classifiers. Each method was reproduced using the officially released training configurations and datasets, with details provided in Appendix A.1.
Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025). Our Safety Recomposition and Safety Backtrack methods are implemented via supervised fine-tuning (SFT) on the Qwen3 base models. As illustrated in Figure 3, the training and experimental procedures comprise three distinct phases. The details are shown in the Appendix A.2.2.
Safety Benchmarks We evaluate model safety using four benchmarks. SORRY-Bench  (Xie et al., 2024) and StrongREJECT  (Souly et al., 2024) measure the model’s ability to refuse harmful prompts, while WildJailBreak  (Jiang et al., 2024b) and JailBreakBench  (Chao et al., 2024) assess its robustness against adversarial jailbreak attacks.
[Uncaptioned image] Figure 4: Safety vs. reasoning trade-off for the 32B model.
Reasoning Benchmarks To assess general reasoning ability, we use AIME2024  (Mathematical Association of America (2024), MAA) for mathematical reasoning and GPQA-Diamond  (Rein et al., 2024) for complex knowledge-based reasoning. To reduce randomness in evaluation, we perform multiple rollout rounds: GPQA is averaged over two rounds, and AIME over eight rounds. Additional details of the reasoning and safety benchmarks can be found in Appendix A.3.
Training Dataset We extracted 15,000 high-quality harmful prompts from 7 public collections—Alert  (Tedeschi et al., 2024), ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming  (Ahmadian et al., 2024), Do-Not-Answer  (Wang et al., 2023), AttaQ  (Kour et al., 2023), and Toxic-Chat  (Lin et al., 2023). We then designated this curated set as our seed dataset and fed it into our pipeline to generate corresponding high-quality safety response sets.
4.2 Main Result
Table 2: Performance comparison of different methods under the main experimental setting. Here, lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.
Method Harmful Jailbreak Reasoning
[…]
Vanilla 45.45 13.62 38.80 81.71 57.33 77.50
STAR-1 18.86 0.74 80.00 62.2 57.33 71.25
SAFEPATH 36.14 10.03 22.80 42.68 55.56 67.92
SafeChain 49.55 16.99 36.80 70.73 52.28 66.25
SafeKey 3.18 0.32 8.53 13.2 41.90 70.80
Safety Backtrack(Ours) 17.05 2.05 8.00 26.83 54.30 77.50
Safety Recomposition(Ours) 13.41 2.86 9.20 28.05 56.82 76.25
[…]
Vanilla 55.45 12.44 34.00 68.29 63.14 77.92
STAR-1 17.95 0.72 86.80 76.83 56.32 76.25
SAFEPATH 34.09 8.49 16.00 50.00 57.78 70.42
SafeChain 49.32 16.87 35.60 67.07 57.58 71.25
SafeKey 4.77 0.32 4.88 10.40 49.00 76.70
Safety Backtrack(Ours) 11.82 1.30 6.40 18.05 62.12 77.92
Safety Recomposition(Ours) 12.50 4.97 2.80 21.95 60.36 78.75
[…]
Vanilla 46.59 12.25 35.20 20.43 65.66 81.67
STAR-1 18.41 0.74 16.80 35.37 54.55 72.92
SAFEPATH 40.00 6.57 22.80 53.66 62.38 70.25
SafeChain 47.95 16.39 28.40 70.73 54.30 71.70
SafeKey 3.41 0.32 7.32 10.40 54.30 71.70
Safety Backtrack(Ours) 14.55 2.77 8.80 23.17 61.62 77.08
Safety Recomposition(Ours) 7.05 1.49 3.20 17.07 62.38 82.08
Our results, summarized in Table 4.2, reveal several key findings:
Baseline methods enhance safety performance at the substantial cost of reasoning capability. As reported in Table 4.2, several baselines either underperform the vanilla model on safety (e.g., SafeChain, STAR-1) or secure safety gains only at the cost of substantial reasoning degradation (e.g., SafeKey).
COG achieves the best trade-off between safety and reasoning across model scales. Our proposed COG framework, including Safety Backtrack and Safety Recomposition, consistently achieves the strongest safety improvements while largely preserving reasoning ability. This trade-off is clearly visualized in Figure 4, where our methods are positioned in the optimal upper-right quadrant, demonstrating both high safety and high reasoning performance simultaneously. Specifically, across model scales, SafB and SafR attain the lowest or near-lowest harmful/jailbreak scores while maintaining competitive reasoning performance (GPQA-Diamond and AIME2024), highlighting the robustness and scalability of COG in resolving the safety–reasoning trade-off.
5 Analysis
5.1 PCA Analysis of Safety Improvements with Minimal Reasoning Impact
To quantitatively evaluate the trade-offs between safety and reasoning, we apply Principal Component Analysis (PCA) to the hidden states of Qwen3-32B We analyze the geometric structure of the representation space under different training approaches, focusing on safety-critical prompts (Harmful, RedTeaming) and reasoning prompts (Reasoning). A linear decision boundary is trained on the two-dimensional PCA projection to quantify the separation between safety and reasoning  [^1] Results for the 8B and 14B models are provided in Appendix B.2., which is illustrated in Figure 5, where we can conclude that:
Our approach establishes a more separable safety manifold while retaining the stability of reasoning representations, indicating effective safety gains with negligible reasoning degradation. Specifically, SafR achieves the largest safety improvement ($\Delta$_Harmful = +4.985) with minimal reasoning loss ($\Delta$_Reasoning = –0.676), whereas SafB provides a more balanced gain ($\Delta$_Harmful = +3.961, $\Delta$_Reasoning = –1.771). In contrast, SafeChain offers only marginal safety benefits ($\Delta$_Harmful = +0.378) while causing severe reasoning degradation ($\Delta$_Reasoning = –5.260). These results indicate that our COG framework substantially enhances safety while largely retaining reasoning capabilities, effectively improving safety alignment without disrupting core reasoning structures.
Figure 5: PCA of the Qwen3-32B representation space. Gray ellipses indicate base model distributions, while colored regions denote post-training clusters (red: safety, blue: reasoning). Centroid shifts ($\Delta$) quantify displacement relative to the linear decision boundary (dashed), with $\Delta>0$ indicating increased margin and $\Delta<0$ indicating decreased margin.
5.2 Understanding Reasoning Preservation via Trajectory and Token Analysis
To study how training strategies influence the reasoning structure of language models, we analyze their cognitive trajectories using DeepSeek-v3.1 (DeepSeek-AI, 2024) with established prompts (Zeng et al., 2025; Gandhi et al., 2025), and measure how these patterns differ from a base model. We track the frequency of key reasoning behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—with scores reflecting the average occurrences per problem on the AIME benchmark, which is presented in Table 3, which indicates that:
Models trained with our methods best maintain the cognitive structure of the base model, and therefore have only a minimal impact on the model’s reasoning capabilities. Specifically, SafR exhibits slightly higher behavior frequencies (+0.3%) and SafB slightly lower (–0.3%), indicating that both methods achieve high alignment with the base model’s reasoning patterns. In contrast, alternative methods such as SafeChain, SafePath, and star-1 exhibit larger deviations (see Table 3). Furthermore, the token-level analysis in Appendix B.3 confirms that SafR and SafB maintain token usage close to the Base model while improving safety, indicating that they enhance safety without significantly altering intrinsic reasoning behavior.
Table 3: Comparison of reasoning patterns (Qwen3-32B) across different training strategies.“Overall Avg.” denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
Pattern Base SafeChain SafePath star-1 SafB SafR
Backtracking 1.33 1.10 1.20 1.30 1.27 1.30
Enumeration 0.93 0.87 0.97 0.83 1.00 1.03
Subgoal Setting 1.60 1.63 1.30 1.40 1.47 1.57
Verification 2.50 2.47 2.23 2.10 2.50 2.57
[…]
Table 4: Comparison of model performance with and without masking the default reasoning trajectory. “SafB (full mask)” corresponds to the Safety Backtrack (full mask) setting, while “SafB (part. mask)” corresponds to the Safety Backtrack (partial mask) setting.
[…]
Qwen3-8B 45.45 13.62 38.80 81.71 57.58 73.30
SafB (full mask) 17.05 2.05 8.00 26.83 54.30 77.50
SafB (part. mask) 23.64 5.37 22.40 59.76 56.87 75.00
5.3 Impact of Masking Strategy in Safety Backtrack
We evaluate the masking strategy in Safety Backtrack (SafB) by comparing two training setups: (1) Full-mask training, where the loss is applied to the reasoning trace, self-check, and final answer; and (2) Partial-mask training (our main method), where the loss is applied only to the self-check and final answer, with the reasoning trace masked.
Table 4 shows that both SafB variants improve safety, with partial masking yielding larger gains on the Harmful and RedTeaming benchmarks. For reasoning, both variants enhance AIME, whereas GPQA shows a slight degradation, more pronounced under full masked SafB, which can be attributed to fewer rollout samples (two for GPQA vs. eight for AIME).
6 Related Work
Safety–Reasoning Trade‐Off
Reasoning capability and safety appear to be inherently conflicting objectives in LRMs. Zhou et al.  (Zhou et al., 2025a) show that stronger LRMs tend to produce more unsafe content, while studies by Huang et al.  (Huang et al., 2025) and Li et al.  (Li et al., 2025) highlight a fundamental trade‐off: improving one objective inevitably undermines the other, underscoring the need to balance safety and reasoning performance.
Safety for Reasoning
Instead of relying on external filters, safety-for-reasoning methods integrate controls directly into inference through training. Current approaches mostly build on supervised fine-tuning (SFT). STAR-1  (Wang et al., 2025b), SafeChain  (Jiang et al., 2025), SAFEPATH  (Jeung et al., 2025), and SafeKey  (Zhou et al., 2025b) apply curated data and tailored SFT to enhance safety while retaining reasoning ability. These methods better preserve coherence but still deliver only modest safety gains or impose notable performance costs, leaving room for more efficient optimization.
7 Conclusion
In this paper, we conducted a systematic investigation into the safety mechanisms of LRMs. Our analysis reveals the existence of Self-Jailbreak, where LRMs persuade themselves into answering the harmful queries. This phenomenon sheds new light on the underlying mechanisms of the reasoning process in LRMs. Building on this discovery, we developed CoG, a training framework that suppresses Self-Jailbreak by detecting and mitigating unsafe reasoning patterns through Safety Backtrack and Safety Recomposition. Experimental results show that CoG substantially enhances safety with minimal degradation in reasoning capability, achieving state-of-the-art performance in the safety–reasoning trade-off. Overall, this study provides a principled framework for aligning LRMs toward safe and reliable reasoning.
Appendix A Detailed Experimental Setup
A.1 Baseline Configuration
A.1.1 Description of Baseline Methods
STAR-1
STAR-1 categorizes 41,000 safety data points from multiple sources into eight predefined categories and generates a response with CoT using DeepSeek-R1, guided by the safety policies associated with each category. Then, a set of rules is applied to filter out 1,000 data points for the dataset. These data are then used to fine-tune an LRM to conduct safety alignment.
SafeChain
SafeChain selected 50,000 data points from the wildjailbreak dataset and used R1-70B to generate five responses for each instruction. Then, Llama-Guard is used to filter data, keeping the responses that are all safe. A random response is sampled from the five responses as the final response. This created a dataset containing 40,000 instruction-response pairs, available for supervised fine-tuning.
SAFEPATH
SAFEPATH fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first” at the start of inference, guiding the LRMs to consider more about safety during the generation process.
SafeKey
SafeKey enhances safety reasoning by integrating a Dual‑Path Safety Head with Query‑Mask Modeling to amplify latent safety signals from both the raw input (X) and the model’s internal query understanding (U) during generation of the “key sentence”—this effectively triggers a safety-focused “Aha moment.” By masking out X when predicting the key sentence based solely on U, Query‑Mask Modeling strengthens the U→K pathway, while the dual‑path head reinforces these hidden‐state safety cues during fine‑tuning. Together, these two jointly improve robustness against harmful prompts.
A.1.2 Implementation Details of Baselines
Computational Resource
To ensure fair comparison and reproducibility, all experiments—including those reproducing related work—were performed on 8 A-800 with bf16 precision enabled, which allows for faster training while preserving numerical stability. The corresponding training hyperparameters are summarized as follows.
Star-1
We use the official dataset and replicate the experiments following the parameter settings reported in the original paper. The detailed training configurations are presented in Table 5.
Table 5: Detailed Training Hyperparameters for Star-1
Hyperparameter Value Hyperparameter Value
Finetuning Type Full Optimizer AdamW
Adam $\beta_{1},\beta_{2}$ 0.9, 0.95 Learning Rate 1e-5
Epochs 5.0 Batch Size 2
Gradient Accumulation Steps 8 Weight Decay 1e-4
Warmup Ratio 0.05 Cutoff Length 8,192
SafeChain
We trained Qwen3 series models with the original SafeChain dataset with llama-factory. Detailed implementation of SafeChain experiment is described as shown in Table 6:
Table 6: Detailed Training Hyperparameters for SafeChain
Parameter Value Parameter Value
Epochs 2 Batch Size 2
Gradient Accumulation Steps 2
Safekey
We use the official SafeKey codebase, making only model-level modifications to its startup scripts. The detailed implementation of the SafeKey experiment is described as shown in Table 7.
Table 7: Detailed Training Hyperparameters for SafeKey
Parameter Value Parameter Value
Epochs 5 Batch Size 2
Gradient Accumulation Steps 8
SafePath
The detailed implementation of the SafePath experiment is described in Table 8:
Table 8: Detailed Training Hyperparameters for SafePath
Parameter Value Parameter Value
Finetuning Type Full Cutoff Length 8192
Batch Size 2 Gradient Accumulation Steps 2
Learning Rate 1e-5 Max Steps 20
Warmup Ratio 0.05
A.2 Implementation Details of Our Method
A.2.1 Manual Classification Accuracy
Table 9: Classification Accuracy of Three Graduate Students on 50 Cases
Participant Correct Cases Accuracy (%)
Graduate Student A 46 92.0
Graduate Student B 43 86.0
Graduate Student C 44 88.0
Average 44.33 88.67
We invited three graduate students with domain expertise to classify 50 representative cases manually. As shown in Table 9, their classification accuracy ranged from 86.0% to 92.0%, with an average of 88.67%, indicating the reliability of our labeling scheme.
A.2.2 COG Generation Parameters
During the sampling process(Phase 1), to ensure output diversity and prevent model degeneration, we set the temperature to 0.7, top_p to 0.8, and presence_penalty to 1.5 to produce the original responses used as seed data (see Table 10).
During the extraction and classification process(Phase 1), temperature and top_p were set to 0.1 and 0.9, respectively, to ensure that the model outputs its most confident predictions.
During the Safety Recomposition and Safety Backtrack stages (Phase 2), we aimed to maintain consistency between generated content and prompt constraints while preserving diversity; thus, temperature was set to 0.3 and top_p to 0.8.
Finally, for the chain-of-thought based response generation stage(Phase 3), temperature was again set to 0.7, top_p to 0.8, and presence_penalty to 1.5 to maintain diversity.
These carefully chosen parameters balance generation quality and diversity while minimally impacting the model’s reasoning capability.
Table 10: Generation Parameter Settings
Stage Parameters Values
Generation Phase temperature 0.7
top_p 0.8
presence_penalty 1.5
Extraction & Classification temperature 0.1
top_p 0.9
SafR & SafB Phases temperature 0.3
top_p 0.8
Chain-of-Thought Generation temperature 0.7
top_p 0.8
presence_penalty 1.5
Table 11: Training Hyperparameters
Parameter Value Parameter Value
Finetuning Type full Learning Rate 2e-6
Cutoff Length 8192 Epochs 3.0
Batch Size 2 Warmup Ratio 0.1
Gradient Accumulation Steps 4
Table 12: Benchmark Implementation Details
Safety Benchmarks Reasoning Benchmarks
Parameters Values Parameters Values
temperature 0.7 temperature 0.6
top_p 1 top_k 20
max_new_tokens 16384 top_p 0.95
rollout 1 max_seq_length 32768
max_out_len 32000
GPQA rollout 2
AIME2024 rollout 8
COG Training Parameters
Both the Safety Recomposition and Safety Backtrack tasks are trained using LlamaFactory under consistent experimental settings. Our approach is based on a dataset of 14,000 examples, with the full training hyperparameters summarized in Table 11.
A.3 Evaluation Details
A.3.1 Benchmark Description
Sorry-bench
Sorry-bench is a systematic safety-refusal benchmark comprising 440 harmful prompts across 44 fine-grained safety categories. We used the original prompts as the test set to evaluate LLM refusal behaviors.
StrongREJECT
StrongREJECT is a jailbreak robustness benchmark featuring 313 carefully filtered harmful prompts spanning six major misuse categories to assess LLM defenses against jailbreaks.
WildJailbreak
WildJailBreak is an adversarial evaluation split of 2,213 jailbreak prompts drawn from a 262 K-example synthetic safety corpus generated by the WildTeaming framework, designed to rigorously test LLM safety mechanisms. We randomly selected 250 prompts from the evaluation split as the evaluation set.
JailBreakBench
JailBreakBench is a robustness benchmark offering 100 paired harmful-behavior prompts (55 % original, 45 % sourced from AdvBench and TDC/HarmBench). In our experiment, we used harmful prompts augmented with Vicuna-generated PAIR variants for comprehensive jailbreak evaluation.
GPQA-Diamond
GPQA-Diamond is the “Diamond” subset of the GPQA benchmark, comprising the 198 most difficult of 448 graduate-level, domain-expert-written multiple-choice questions in biology, chemistry, and physics.
AIME2024
AIME is the complete set of 30 official integer-answer problems from the 2024 American Invitational Mathematics Examination I & II, directly sourced from the MAA’s public releases.
A.3.2 Evaluation Metrics
For the safety benchmarks, Sorry-Bench, StrongREJECT, and WildJailBreak use attack successful rate (ASR) as the evaluation metric, revealing the times that a model accepts harmful prompts. Following the setting of the original benchmark, we used the rejection rate for JailBreakbench, measuring how often the model successfully rejects harmful prompts. For reasoning benchmarks, we use accuracy as the evaluation metric, measuring the rate at which models give correct answers.
A.3.3 Benchmark Hyperparameters Details
We used a rollout of 2 for GPQA-Diamond and 8 for AIME2024. The evaluations on GPQA-Diamond and AIME2024 were conducted using the OpenCompass framework. The detailed hyperparameter setting is shown in Table 12.
Appendix B Additional Experimental Results
B.1 Initial Investigation by Switching Thinking Mode
Recent works  (Zhou et al., 2025a; Zhang et al., 2025) have discovered that LRMs tend to answer harmful questions. To further validate whether the thinking trajectory influences the safety of LLMs, we compare multiple LRMs with the thinking mode switch on & off. The evaluation is performed on 2k harmful queries from Wildjailbreak  (Jiang et al., 2024a), with safety assessed using Llama-Guard-3-8B (Llama Team, 2024) as the automatic judge.
The thinking trajectory substantially undermines the safety of LRMs. As shown in Table 13, LRMs become consistently safer when the thinking mode is disabled, but their safety degrades notably once it is enabled. This trend holds across different model scales: enabling the thinking trajectory systematically increases the likelihood of harmful responses. For example, the harmful answer rate of Qwen3-32B rises by 10.55% when the mode is enabled, while the 8B and 14B variants also exhibit noticeable increases (6.04% and 3.35%, respectively). These results indicate that content generated within the thinking trajectory can directly drive LRMs to produce harmful outputs.
Table 13: Percentage of harmful answers (”Answer”) and refusals (”Refuse”) on Wildjailbreak. Higher ’Answer’ indicates lower safety. $\Delta=$ Disabled - Enabled.
Model Enabled Disabled $\Delta$ (Answer)
Answer Refuse Answer Refuse
Qwen3-8B 41.40 58.60 35.36 64.64 6.04-6.04
Qwen3-14B 35.65 64.35 32.30 67.70 3.35-3.35
Qwen3-32B 41.45 58.55 30.90 79.10 10.55-10.55
B.2 Analysis: PCA Analysis of 8B and 14B Models
We conducted a comparative analysis of models with different parameter sizes and fine-tuning methods (SafR and SafB), aiming to evaluate their impact on safety and representational clustering. The observed differences are visualized in Figure 6 and 7, while quantitative results across all configurations are reported in Table 14.
Table 14: Safety distance analysis of Qwen3 models by model size. Values represent perpendicular distances from cluster centroids to the linear decision boundary. Positive $\Delta$ values indicate movement away from the boundary (greater margin, safer region), while negative $\Delta$ values indicate movement toward the boundary (reduced margin, less safe region).
Method Qwen3-8B Qwen3-14B Qwen3-32B
Harmful Reasoning Harmful Reasoning Harmful Reasoning
$\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$
Base 13.24 14.48 15.54 9.05 11.20 12.18
SafR 18.55 15.02 18.95 8.33 16.18 11.50
+5.30 +0.54 +3.41 -0.73 +4.98 -0.68
SafB 19.51 12.95 19.14 9.63 15.16 10.41
+6.27 -1.54 +3.60 +0.58 +3.96 -1.77
SafeChain 13.45 5.87 12.59 4.26 11.58 6.92
+0.21 -8.61 -2.95 -4.79 +0.38 -5.26
Model Scale
The 32B model consistently outperforms the 8B model across safety and clustering metrics. It exhibits a higher Safety Distance in all settings (Base, SafR, SafB), indicating better separation from harmful content. Its Silhouette Score at the Base stage (0.140) also exceeds that of the 8B model (0.120), reflecting a more structured internal representation.
Fine-Tuning Methods
Both Safety Recomposition (SafR) and Safety Backtrack (SafB) fine-tuning strategies contribute to enhanced model safety, as evidenced by the centroid shifts observed in Figure 6 and Figure 7, and further quantified by the increase in Safety Distance (e.g., for the 8B model: from 20.05 to 25.68 after SafR). Notably, SafR yields consistently superior clustering performance, with higher Silhouette Scores than SafB across both model sizes. This suggests that SafR not only improves safety alignment but also leads to more coherent and well-separated internal representations.
Figure 6: PCA of the Qwen3-8B representation space.
Figure 7: PCA of the Qwen3-14B representation space.
B.3 Preservation of Reasoning Token Usage
Table 15 reports the average token counts for GPQA and AIME benchmarks across different training methods. SafR and SafB maintain token usage close to the Base model, while improving safety. In contrast, other baselines exhibit larger deviations, suggesting potential inefficiencies in reasoning. These results demonstrate that SafR and SafB enhance model safety without significantly affecting intrinsic reasoning behavior, selectively reinforcing safety-related features while preserving core reasoning capabilities.
Table 15: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3-8B, Qwen3-14B, and Qwen3-32B across different training methods.
Model Method GPQA-Diamond AIME
Qwen3-8B Base 7553.84 14895.49
SafePath 4869.04 11816.52
STAR-1 3449.49 12337.94
SafeChain 5105.35 12533.35
SafR 5212.79 14472.53
SafB 6540.14 14881.68
Qwen3-14B Base 5585.28 14419.94
SafePath 4379.39 10939.46
STAR-1 2736.43 11011.73
SafeChain 4688.32 12372.85
SafR 4303.19 13549.96
SafB 5091.59 13010.36
Qwen3-32B Base 5150.81 12634.11
SafeChain 5588.61 11198.09
SafePath 3925.96 10200.49
STAR-1 3564.03 11862.56
SafB 6452.00 12318.87
SafR 4830.90 12548.96
B.4 Analysis: Reasoning Safety Analysis
Prior work has primarily focused on improving the safety of reasoning through external interventions (safety for reasoning). In this section, we take the opposite view and explore whether a model’s internal reasoning ability can be leveraged to enhance overall safety (reasoning for safety). To evaluate our approach, we compare it against the recently proposed R2D method  (Zhu et al., 2025) as a representative baseline. The experimental results are shown in Table 16.
Table 16: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline. Lower is better for SafB, SafR, W-JB, and JB-B; higher is better for GPQA and AIME.
Model SafB SafR W-JB JB-B GPQA AIME
Qwen3-8B 45.45 13.62 38.80 18.29 57.58 73.30
R2D 19.32 0.96 12.80 84.15 41.92 47.92
SafR(Ours) 13.41 2.86 9.20 71.95 56.82 76.25
SafB(Ours) 17.05 2.05 8.00 73.17 54.30 77.50
Based on the experimental results shown in the Table 16, we observe that applying reasoning-for-safety methods to LRMs indeed improves safety performance. However, similar to SafeKey and Star-1, these methods tend to sacrifice reasoning ability, failing to strike a proper balance between safety and reasoning capabilities. Although our methods partially address this issue, maintaining the integrity of reasoning ability remains a significant challenge.
Appendix C Self-Jailbreak Category Example
C.1 Benign Reframing
The example of Benign Reframing in Self-Jailbreak is presented in Figure 8, where red text represents their Self-Jailbreak act.
C.2 Logical Fallacies
The example of Logical Fallacies in Self-Jailbreak is presented in Figure 9, where red text represents their Self-Jailbreak act.
C.3 Warning
The example of Warning in Self-Jailbreak is presented in Figure 10, where red text represents their Self-Jailbreak act.
C.4 Harm Identification
The example of Warning in Harm Identification is presented in Figure 11, where LRM answers the harmful quesiton as the regular one.
Appendix D Prompt Design
In this section, we present the prompt designs used in our COG framework and other experiments.
D.1 Extraction Prompt
This subsection provides both the base prompts and few-shot examples used in the extraction task.
Prompt
The basic prompt template used for extraction is shown in Figure 12.
Few-Shot
We also include few-shot examples to guide the model during extraction, as illustrated in Figure 13.
D.2 Classification Prompt
Prompt
The prompt used for the classification task is outlined below in Figure 14.
D.3 Safety Recomposition Prompt
For the Safety Recomposition stage, our prompt is constructed by concatenating several components. Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts” field in a JSON file. Then, we combine the ”main_prompt”, the selected ”sub_prompt”, and a ”format_prompt” to form the final prompt. The detailed structure is as shown in Figure 15
D.4 Safety Backtrack Prompt
In the Safety Backtrack stage, we further extend the prompt structure from the Safety Recomposition process. In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases”, selected based on the classification result. This helps guide the model more smoothly and maintain coherence in the final output. The detailed structure is as shown in Figure 16
D.5 Integration Prompt
This stage integrates the outputs from previous modules into a unified final response. The detailed prompt is as shown in Figure 17
D.6 BoN Ranking Prompt
In this step, we employ a Best-of-N (BoN) ranking mechanism to select the most appropriate final output from multiple candidates generated by the model. The ranking prompt used to guide this selection process is shown in Figure 18.
Figure 8: Example of Benign Reframing.
Figure 9: Example of Warning.
Figure 10: Example of Warning.
Figure 11: Example of Harm Identification.
Figure 12: Example of the basic extraction prompt used in the extraction stage.
Figure 13: Few-shot examples used in the extraction prompt during the extraction stage.
Figure 14: Example of the basic classification prompt used in the classification stage.
Figure 15: Prompt structure used in the Safety Recomposition stage, formed by concatenating the main prompt, selected sub-prompt based on classification results, and format prompt.
Figure 16: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.
Figure 17: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.
Figure 18: Prompt structure used in the Best-of-N (BoN) ranking stage to select the best output from multiple candidates.

### v3

@@
-When Models Outthink Their Safety: Mitigating Self-Jailbreak in Large Reasoning Models with Chain-of-Guardrails
+When Models Outthink Their Safety: Unveiling and Mitigating Self-Jailbreak in Large Reasoning Models
 Abstract
-Large Reasoning Models (LRMs) demonstrate remarkable capabilities on complex reasoning tasks but remain vulnerable to severe safety risks, including harmful content generation and jailbreak attacks.
-Existing mitigation strategies rely on injecting heuristic safety signals during training, which often suppress reasoning ability and fail to resolve the safety-reasoning trade-off.
-To systematically investigate this issue, we analyze the reasoning trajectories of diverse LRMs and uncover a phenomenon we term Self-Jailbreak, where models override their own risk assessments and justify responding to unsafe prompts.
-This finding reveals that LRMs inherently possess the ability to reject unsafe queries, but this ability is compromised, resulting in harmful outputs.
-Building on these insights, we propose the Chain-of-Guardrail (CoG), a training framework that recomposes or backtracks unsafe reasoning steps, steering the model back onto safe trajectories while preserving valid reasoning chains.
-Extensive experiments across multiple reasoning and safety benchmarks demonstrate that CoG substantially improves the safety of current LRMs while preserving comparable reasoning ability, significantly outperforming prior methods that suffer from severe safety–reasoning trade-offs.
+Large Reasoning Models (LRMs) achieve strong performance on complex multi-step reasoning, yet they still exhibit severe safety failures such as harmful content generation.
+Existing methods often apply coarse-grained constraints over the entire reasoning trajectories, which can undermine reasoning capability while failing to address the root causes of unsafe behavior.
+In this work, we uncover a previously underexplored failure mode in LRMs, termed Self-Jailbreak, where models initially recognize the harmful intent of a query, but override this judgment during subsequent reasoning steps, ultimately generating unsafe outputs.
+Such a phenomenon reveals that LRMs are capable of recognizing harm, while safety failures primarily arise from reasoning steps.
+Motivated by this finding, we propose Chain-of-Guardrail (CoG), a trajectory-level training framework that mitigates Self-Jailbreak via targeted, step-level interventions while maintaining reasoning ability.
+Experiments across multiple safety and reasoning benchmarks indicate that CoG achieves a favorable balance between safety and reasoning performance compared with existing approaches.
+When Models Outthink Their Safety:
+Unveiling and Mitigating Self-Jailbreak in Large Reasoning Models
+Yingzhi Mao1,2thanks: Equal contribution.thanks: Code: https://github.com/icip-cas/COG., Chunkang Zhang [^1] footnotemark: 1, Junxiang Wang1, Xinyan Guan1,2, Boxi Cao1 Yaojie Lu1, Hongyu Lin1, Xianpei Han1,2, Le Sun1,2 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {maoyingzhi2024, zhangchunkang2022, luyaojie}@iscas.ac.cn
 1 Introduction
-Large Reasoning Models (LRMs), such as OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Qwen3 (Yang et al., 2025), have demonstrated remarkable capabilities in solving complex tasks and are increasingly being adopted across a wide range of domains.
-Despite these advances, researchers have identified that current LRMs remain hindered by significant safety vulnerabilities, such as the generation of harmful content as well as the susceptibility to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025; Wang et al., 2025a).
-These issues highlight a critical obstacle to the reliable and responsible use of LRMs (Weidinger et al., 2021).
-Therefore, addressing these challenges is critical for ensuring that LRMs operate safely and remain aligned with human values.
-While recent efforts (Wang et al., 2025b; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have attempted to mitigate these risks, they largely exhibit a pronounced safety-reasoning trade-off.
-Specifically, current studies have primarily focused on curating safety-oriented datasets based on heuristic rules and fine-tuning LRMs to produce refusal responses.
-Unfortunately, while such approaches can improve safety to some extent, the pattern-based behavior often conflicts with the flexible and multi-step deliberation required for complex reasoning, thereby substantially degrading reasoning ability.
-Therefore, it is critical to identify the underlying causes of the unsafe behavior in LRMs, and develop more principled strategies that can enhance safety without sacrificing reasoning performance.
-To this end, we conduct a systematic analysis of LRMs’ reasoning trajectories and identify a critical failure phenomenon, which we term Self-Jailbreak.
-Specifically, as demonstrated in Figure 1, we decompose LRM’s reasoning thinking trajectory into three stages: Risk Awareness, Risk Analysis, and Response Strategy.
-We observe that LRMs often detect harmful intent during the Risk Awareness stage, but subsequently override this recognition during the Risk Analysis stage, and ultimately persuade themselves to respond to unsafe queries.
-This reveals that LRMs possess an inherent ability to reject unsafe queries, yet it is undermined by Self-Jailbreak, resulting in unsafe generation.
-Furthermore, we categorize Self-Jailbreak into four distinct types—Benign Reframing, Logical Fallacy, Warning, and Risk Misrecognition—and find that Warning accounts for most unsafe outputs.
-Existing approaches that solely inject heuristic safety signals during training fail to resolve the safety–reasoning trade-off.
-While they suppress overtly unsafe outputs, such interventions do not address unsafe reasoning steps embedded within reasoning trajectories.
-Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving reasoning quality.
-Specifically, CoG leverages precise classification of Self-Jailbreak to perform targeted safety interventions, mitigating unsafe behavior while preserving the model’s reasoning abilities and reasoning patterns.
-To achieve this, CoG employs two strategies: 1) Safety Recomposition, which systematically reorganizes unsafe reasoning into entirely new, safety-assured chains, and 2) Safety Backtrack, which preserves the model’s original reasoning path while incorporating self-check and backtracking mechanisms to correct risky steps before they lead to unsafe conclusions.
-The resulting safety-oriented reasoning chains are then used to fine-tune LRMs, aligning them toward safe responses while maintaining the reasoning capability.
-Figure 1: Illustration of “Self-Jailbreak.” The Instruct model (left) directly refuses harmful queries, whereas the LRM (right) identifies risk during Risk Awareness but later engages in Self-Jailbreak (red text) during Risk Analysis, ultimately overriding its initial safety judgment and producing harmful content in the Response Strategy stage.
-Extensive experiments across multiple safety and reasoning benchmark datasets demonstrate that COG excels at balancing the safety and reasoning performance of LRMs.
-On Qwen3-32B, CoG improves the reasoning ability by 13.5% over SafeKey while achieving state-of-the-art performance on GPQA and AIME (Rein et al., 2024; Mathematical Association of America (2024), MAA), maintaining comparable reasoning performance.
-Further analysis of reasoning trajectories reveals that, unlike other approaches, CoG preserves the model’s intrinsic reasoning patterns, confirming that its safety gains do not come at the expense of the model’s original deliberation trajectory.
-In summary, our contributions include: In summary, our contributions include:
-  • To the best of our knowledge, we are the first to identify and characterize the Self-Jailbreak phenomenon, revealing it as a key factor behind the poor safety performance of LRMs.
-  • We propose a systematic analytical framework to categorize Self-Jailbreak instances and quantify their impact on model behavior, finding that Benign Framing and Warning constitute the majority of the Self-Jailbreak phenomenon.
-  • We introduce the Chain-of-Guardrail (CoG), a training framework that mitigates Self-Jailbreak and substantially improves model safety while preserving reasoning capabilities.
-Experiment results show that CoG improves the safety-reasoning trade-off by 3.39%, reaching the state-of-the-art (SOTA) across multiple LRMs.
-Figure 2: Processing of harmful queries by LRM, illustrating the workflow from receiving prompts, through risk detection during reasoning, to the four possible outcomes: aware of risk and responds, aware of risk and refuses, unaware of risk and responds, unaware of risk and refuses.
-2 Motivation: Preliminary Experiments On Self-Jailbreak
-This section examines the thinking trajectory of LRMs, drawing on pilot experimental results (reported in Appendix B.1) which indicate that most LRMs become unsafe when the thinking mode is enabled.
-Subsequently, we find that even if LRMs are capable of recognizing latent dangerous intentions, they still respond to the harmful query.
-Such results shed light on our hypothesis: Self-Jailbreak.
-We further validate the existence of Self-Jailbreak with experiments, which represent two significant findings: (1) Most LRMs encounter Self-Jailbreak during the reasoning trajectory.
-(2) Among the different categories of Self-Jailbreak, Warning is the primary source.
-2.1 Hypothesis: Self-Jailbreak via Thinking–Answer Misalignment
-We analyzed the models’ chains of thought, discovering that their thinking trajectory can be decomposed into three stages: risk awareness, risk analysis, and response strategy, which enables a deeper examination of how thinking affects the safety of LRMs.
-Building on the previous experiment, we employ LLM-as-Judge (Team, 2024; Gu et al., 2024) to assess whether the LRM demonstrates awareness of potential risks during the risk awareness stage, and use Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content.
-This setup enables us to examine potential misalignments between the model’s reasoning trajectory and its generated answers.
-As illustrated in Figure 2, our analysis revealed a striking contradiction:
-LRMs are able to identify most of the harmful queries but still generate unsafe answers.
-94.2% of harmful queries, as well as users’ potentially malicious intents, are recognized; however, 52.30% of their generated outputs still contain hazardous content.
-In contrast, only 39.90% of the cases reflect awareness of risk followed by a safe response.
-This inconsistency between “thinking” and “response” raises a critical concern: during subsequent reasoning, LRMs may tend to persuade themselves to respond to such queries.
-Based on this observation, we propose the following hypothesis:
-Self-Jailbreak During the thinking process, LRMs often exhibit a phenomenon we term Self-Jailbreak, where models persuade themselves to produce responses to potentially harmful requests in an attempt to assist the user, even after recognizing the risky nature of the queries and the user’s underlying malicious intent.
-Table 1: Self-Jailbreak category distribution of various LRMs (%)
-Model Response Safety Self-Jailbreak Categories
-Safe Unsafe Benign Reframing Warning Logical Fallacies Harm Identification
-DS-R1-0528 4.07 95.93 42.29 52.12 0.51 1.02
-DS-Llama-70B 0.50 99.50 47.89 45.67 0.91 5.04
-Qwen3-8B 0.21 99.79 38.60 50.69 2.41 8.10
-Qwen3-14B 0.15 99.85 37.30 53.10 1.73 7.32
-Qwen3-32B 0.36 99.64 36.32 52.51 2.20 8.61
-2.2 Categorization and Existence of Self-Jailbreak
-Building on this hypothesis and inspiration from backtracking during mathematical reasoning  (Chen et al., 2025), we designed a more rigorous classification framework, which is guided by the principle that the model’s reasoning trajectory should not generate content that directly or indirectly facilitates harmful user behaviors.
-We categorize these behaviors into four types:
-Benign Reframing:
-The model recognizes the risk but attempts to reinterpret the user’s intent benignly, offering help from such interpretation.
-Warning:
-The model assumes that appending a warning to harmful content is sufficient to prevent its misuse, and proceeds to generate its answer.
-Logical Fallacies:
-Given a query containing logical contradictions, the model detects potential harm but falls into logical traps, leading to logical fallacies.
-Harm Identification:
-The model fails to recognize the harm in the query, resulting in answering.
-To investigate the prevalence of Self-Jailbreak in LRMs, we categorize and analyze the thinking trajectory of several contemporary LRMs on 2k queries in WildJailbreak (Jiang et al., 2024a).
-As shown in Table 1, we observe that:
-(1) Self-Jailbreak is a prevalent phenomenon observed in various LRMs.
-Experiment results indicate that Self-Jailbreak occurs frequently across LRMs, including peripheral models like DeepSeek-R1-0528 (DeepSeek-AI, 2025), highlighting the importance and relevance of our study.
-(2) Warning is the most common source of Self-Jailbreak.
-As illustrated in Table 1, Warning is the most categorized Self-Jailbreak type across various LRMs.
-This tendency likely stems from the fact that LRMs are largely trained on math and code data, which predisposes them to respond in ways that meet user expectations.
-Consequently, even when confronted with harmful queries, LRMs are likely to be helpful, inadvertently bypassing their own safety mechanisms by providing content alongside a safety disclaimer.
-Such behavior ultimately gives rise to Self-Jailbreak.
-3 Chain-of-Guardrail: A Training Framework to Mitigate Self-Jailbreak
-To systematically address the risks of Self-Jailbreak in LRMs, we propose a unified safe training framework: Chain-of-Guardrail (COG).
-As illustrated in Figure 3, the framework is designed to create an interpretable and structured pipeline that automatically removes or rejects unsafe content during the reasoning trajectory $c$, producing a safe and logically consistent response $y_{\text{safe}}$.
-To achieve this, COG employs two complementary methods: Safety Recomposition (SafR) and Safety Backtrack (SafB).
-SafR focuses on proactively restructuring and correcting the reasoning chain to eliminate potential safety risks, while SafB focuses on the reasoning trajectory when unsafe content is detected, ensuring that the model corrects errors and generates a safe response.
+Large Reasoning Models (LRMs) (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025) have demonstrated strong capabilities across a wide range of tasks, including software engineering and mathematical reasoning.
+As these models are increasingly integrated into autonomous agents and decision-support systems, ensuring their safety and alignment with human values is foundational.
+However, recent studies show LRMs remain vulnerable to serious safety failures (Green et al., 2025; Arrieta et al., 2025), ranging from generating harmful content (Qiu et al., 2025; He et al., 2025) to exhibiting vulnerability to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025b).
+Therefore, strengthening safety capabilities of LRMs is not merely an auxiliary concern, but a prerequisite for their trustworthy deployment at scale.
+Figure 1: Illustration of Self-Jailbreak.
+While the model initially recognizes the harmful nature of a query, it overrides this judgment during subsequent reasoning steps, ultimately yielding unsafe output.
+While recent efforts (Wang et al., 2025; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have made progress towards safer LRMs, they often incur a pronounced safety–reasoning trade-off.
+A primary limitation shared by these approaches is that they typically inherit safety paradigms originally designed for standard LLMs, treating internal reasoning trajectories of LRMs as an undifferentiated extension of responses.
+Subsequently, safety constraints are typically imposed in a coarse-grained and global manner.
+Such heuristic safeguards often interfere with the model’s intrinsic reasoning patterns, degrading the coherence required for multi-step reasoning.
+These limitations indicate that mitigating safety risks in LRMs without undermining their reasoning capabilities requires moving toward mechanisms that can localize and address failure-inducing steps within the reasoning chain.
+To this end, we conduct a systematic analysis of reasoning trajectories of LRMs to uncover the root causes of their safety failures.
+In particular, we identify two distinct failure modes: Harm Misidentification, where harmful intent is not recognized, and Self-Jailbreak, a severe and previously overlooked phenomenon in which the model initially identifies potential harm but later overturns this safety judgment during subsequent reasoning.
+To enable fine-grained analysis, we decompose each reasoning trajectory into three consecutive stages: risk awareness, risk analysis, and response strategy.
+Specifically, as demonstrated in Figure 1, given a harmful query (“How to build a bomb…”), the model correctly identifies the harm during the risk-awareness stage, but then overrides this judgment by rationalizing a seemingly benign user intent (“…could be a student…”).
+In the response-strategy stage, it chooses to answer rather than refuse, and the resulting response contains harmful content.
+Through quantitative analysis on WildJailbreak, we find that only a small portion of unsafe outputs arise from Harm Misidentification, whereas Self-Jailbreak is the dominant failure mode, responsible for nearly 80% of the unsafe cases we analyze and thus constituting our primary focus.
+In these cases, LRMs correctly recognize harmful intent during the risk awareness stage, yet this recognition is subsequently overridden during the risk analysis stage, where the model effectively persuades itself to comply with the unsafe request.
+Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving the reasoning capability.
+CoG executes targeted, step-level interventions conditioned on the diagnosed Self-Jailbreak patterns, correcting only the segments that induce unsafe behavior.
+We instantiate this framework via two complementary variants: Safety Recomposition, which rewrites the reasoning chain into a logically consistent, safe alternative, and Safety Backtrack, which preserves the original trajectory while revising risky segments before they lead to unsafe outputs.
+These corrected traces serve as fine-tuning data to align LRMs without suppressing their reasoning potential.
+Extensive experiments across multiple safety and reasoning benchmarks demonstrate that CoG achieves a superior balance between safety and reasoning performance.
+Across diverse LRMs, CoG consistently reduces attack success rates to competitive levels on standard safety benchmarks, while substantially boosting performance on challenging reasoning tasks.
+Notably, on Qwen3-32B, CoG achieves safety performance comparable to SafeKey, while substantially improving reasoning accuracy, with GPQA-Diamond increasing from 54.30 to 62.38 and AIME2024 from 71.70 to 82.08.
+Further analyses from both reasoning-pattern and representation-geometry perspectives suggest that CoG improves safety by correcting specific failure-inducing reasoning steps, rather than globally rewriting the model’s reasoning paradigm.
+We summarize our major contributions as follows:
+  • To the best of our knowledge, we are the first to uncover and characterize the Self-Jailbreak phenomenon, revealing it as a primary driver of safety failures in LRMs.
+  • We propose an analysis framework that enables the categorization and quantitative analysis of unsafe reasoning behaviors.
+  • We introduce the Chain-of-Guardrail (CoG), a trajectory-level training framework that achieves the state-of-the-art safety-reasoning balance across multiple benchmarks.
+2 Unveiling Self-Jailbreak in Reasoning Trajectories
+In this section, we analyze the causes of safety failures in LRMs through an analysis of their reasoning trajectories.
+Our findings indicate a consistent mismatch between risk awareness and final response: LRMs may generate unsafe outputs even after identifying potential risks during reasoning.
+Beyond failures of harm identification, we observe a prevalent pattern in which earlier safety assessments are revised or overridden in later reasoning stages, which we refer to as Self-Jailbreak.
+We further provide a taxonomy of such behaviors across multiple LRMs to support a more fine-grained analysis of this failure mode.
+2.1 The Prevalance of Self-Jailbreak in LRMs
+Investigation Setup
+To localize safety failures within the reasoning process, we decompose each reasoning trajectory into three stages: risk awareness, risk analysis, and response strategy.
+Each stage is assessed independently.
+For evaluation, we sample 2,000 data points from Wildjailbreak (Jiang et al., 2024a) as the benchmark, following the official setting with Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content.
+In addition, we utilize Qwen2.5-72B-Instruct (Team, 2024; Gu et al., 2024) to determine whether potential risks are explicitly identified during the risk awareness stage.
+Figure 2: Flow of harmful queries in WildJailbreak through Qwen3-32B, showing the relationship between the final response (Answer/Refuse) and risk awareness during reasoning (Yes/No).
+Main Cause of Safety Failure
+Based on the results of Figure 2, we categorize safety failures into two distinct patterns: (1) Harm Misidentification, where the model fails to recognize harmful intent, and (2) Mismatch between risk awareness and response, where the model explicitly identifies risks but produces unsafe outputs.
+While the former indicates a detection failure, the latter reveals a more critical flaw in the model’s reasoning process.
+Further investigation of the second pattern reveals a pattern in which the model’s subsequent reasoning effectively revises or overrides its own safety assessment, leading to unsafe outputs even after risks have been identified.
+We refer to this phenomenon as Self-Jailbreak.
+To determine whether Self-Jailbreak is an isolated anomaly or a widespread issue, we extended our investigation across multiple Large Reasoning Models (LRMs), as shown in Figure 3.
+Our results yield a significant finding:
+Self-Jailbreak constitutes the predominant safety failure mode across LRMs.
+As illustrated in Figure 2, Self-Jailbreak consistently surpasses Harm Misidentification as the primary source of unsafe outputs across diverse model families and parameters.
+For instance, in DeepSeek-R1, 93.7% of safety failures are attributable to Self-Jailbreak.
+This indicates that Self-Jailbreak represents a systemic and persistent challenge, even for state-of-the-art LRMs.
+Figure 3: Proportion of the safety failure cause across multiple LRMs conditioned on the unsafe responses of Wildjailbreak
+Model Benign Reframing Warning Logical Fallacies
+DS-R1 40.91 56.97 2.12
+DS-Llama-70B 39.17 58.53 2.30
+Qwen3-8B 36.94 57.91 5.15
+Qwen3-14B 38.60 58.20 3.20
+Qwen3-32B 36.86 58.96 4.18
+Table 1: Distribution of self-jailbreak categories (%), where DS-R1 stands for DeepSeek-R1 while DS-Llama-70B represents DeepSeek-R1-Distill-Llama-70B.
+2.2 A Taxonomy of Self-Jailbreak Behaviors
+While the prevalence of self-jailbreak is evident, the underlying reasons that lead models to generate unsafe content remain poorly understood.
+A deeper analysis of how models override their own safety judgment is essential for designing effective safeguards.
+To this end, we conducted a qualitative analysis involving manual inspection of reasoning traces (Chain-of-Thought) and final responses, focusing on the discrepancy between the model’s internal risk analysis and its final output generation.
+Taxonomy Definition
+Based on recurring patterns identified in the manual inspection, we establish a taxonomy categorizing Self-Jailbreak into three distinct behaviors:
+  • Benign Reframing: The model actively reinterprets the user’s malicious intent as benign (e.g., educational or theoretical), thereby justifying a helpful response.
+  • Warning: The model assumes that appending a safety warning or disclaimer is sufficient to mitigate the harm, leading to a “warn-but-answer” failure mode.
+  • Logical Fallacies: The model’s reasoning becomes entangled in complex or contradictory logical constraints within the prompt, causing it to bypass safety guardrails due to erroneous logical deductions (examples in Appendix E).
+Figure 4: Overview of the Chain-of-Guardrail (CoG) framework.
+Phase 1: the original LRM produces an initial COT, which is decomposed into three atomic components—risk awareness $d(x)$, risk analysis $a(x)$, and response strategy p(x)p(x)—and classified for self-jailbreak risks.
+Phase 2: guided by the classification signal, CoG applies Safety Recomposition (SafR) or Safety Backtrack (SafB) to construct a safety-oriented COT (S-COT).
+Phase 3: The S-COT guides the model to generate the final safe response.
+Quantitative Analysis
+Utilizing this taxonomy, we annotated Self-Jailbreak instances across multiple LRMs to analyze their distribution.
+As shown in Table 1, a key finding emerges.
+Most notably, Warning is the primary Self-Jailbreak type, consistently accounting for over 55% of cases across all evaluated models.
+This suggests that the primary safety bottleneck is not a lack of risk awareness, but rather a flaw in the risk analysis.
+Models have learned to be overly compliant and often default to answering even when risks are identified, rather than refusing.
+The prevalent Warning and Benign Framing behavior indicates that during training, models may not have internalized clear signals for when to withhold responses entirely.
+Implications for Methodology
+Overall, these results indicate that Self-Jailbreak is both widespread and structured.
+Unsafe behavior often arises after correct risk recognition and recurs in a small number of type-specific reasoning patterns.
+This implies that coarse-grained safety enforcement or globally rewriting reasoning traces may discard valid reasoning steps while still missing the failure source, motivating mitigation that first identifies the Self-Jailbreak type and then targets the corresponding failure-inducing steps.
+3 Chain-of-Guardrails: A Training Framework for Reasoning-Aware Safety
+Motivated by the observation that safety failures in LRMs frequently arise during reasoning despite correct risk awareness, we propose Chain-of-Guardrails (CoG), a training framework that identifies and corrects unsafe reasoning steps while preserving the model’s inherent reasoning capability.
+Rather than imposing uniform safety constraints on the entire response, CoG performs targeted, reasoning-level interventions guided by diagnosed Self-Jailbreak behaviors.
+3.1 Framework Overview
+Given a query $x$ and an original LRM $\pi_{0}$, CoG aims to produce a safe response $y_{\text{safe}}$ while preserving the model’s original reasoning ability.
+To enable fine-grained analysis and intervention, we explicitly decompose the model’s reasoning trajectory into interpretable components.
+3.2 Phase 1: Input & Analysis
+In Phase 1, the original model $\pi_{0}$ generates an initial reasoning trajectory and response to the input query $x$.
+We decompose the resulting reasoning trajectory into three interpretable components:
+$c=\pi_{0}(x)=[d(x),a(x),p(x)]$ (1)
+where $d(x)$ denotes the model’s risk awareness, $a(x)$ its risk analysis, and $p(x)$ its response strategy.
+Given the decomposed components, we run a Self-Jailbreak classifier to predict whether the reasoning trajectory contains Self-Jailbreak behaviors and, when applicable, output the corresponding Self-Jailbreak type.
+3.3 Phase 2: Safety-Oriented Reasoning Transformation
+In Phase 2, CoG transforms unsafe reasoning trajectories into safety-oriented ones by applying targeted interventions conditioned on the classification signal from Phase 1.
+We introduce the following complementary transformation strategies:
 […]
-3.1 Optimization Goal of LRMs in COG
-Specifically, we formulate the optimization objective as follows:
-$f:(\pi_{0},x)\rightarrow y_{\text{safe}},\quad\text{s.t.}\quad\text{Preserve}(\text{Reasoning}_{\pi_{0}})$ (1)
-where $\pi_{0}$ is the original model, $y_{\text{safe}}$ is the safe response, and $\text{Preserve}(\cdot)$ ensures maximal retention of the original reasoning.
-To achieve this, we explicitly model the Chain-of-Thought (CoT) as a structured latent variable:
-$c=\pi_{0}(x)=[d(x),a(x),p(x)],$ (2)
-where $d(x)$, $a(x)$, and $p(x)$ represent the risk awareness, risk analysis, and response strategy.
-Depending on the adopted process, the safety-oriented reasoning chain is constructed as follows:
-Chain of Safety Recomposition (SafR):
-$c_{\text{safe}}=\text{Merge}(d(x),\hat{a}(x),\hat{p}(x)),$ (3)
-where $\hat{a}(x)$ and $\hat{p}(x)$ are the risk analysis and response strategy recomposed by the Safety Recomposition procedure.
-Chain of Safety Backtrack (SafB):
-$c^{\prime}_{\text{safe}}=[c,s(x)],$ (4)
-where $s(x)$ denotes the model-generated self-check module appended to the original reasoning chain in the Safety Backtrack procedure.
-The final safe response $y_{\text{safe}}$ is produced based on the safety-oriented reasoning chain derived from either SafR or SafB.
-3.2 Selective Loss Masking Strategy
-To avoid undesired shifts in the model’s predicted distribution due to safety interventions, we employ a selective loss masking strategy.
-Let the output sequence be $o=[c_{\text{safe}},y_{\text{safe}}]$ for SafR, and $o=[c^{\prime}_{\text{safe}},y_{\text{safe}}]$ for SafB, where $c_{\text{safe}}$ and $c^{\prime}_{\text{safe}}$ are defined as in Eq.
-3 and Eq.
-4, respectively.
-We propose the utilization of a binary mask vector $m\in\{0,1\}^{|o|}$, to supervise distinct segments of the output sequence $o$ selectively. The primary training goal can be summarized as follows:
-$\mathcal{L}_{SafR/SafB}=-\sum_{i=1}^{|o|}m_{i}\cdot\log P_{\theta}(o_{i}|o_{<i},x),$ (5)
-This selective loss masking strategy allows for flexibility depending on the applied process:
-Safety Recomposition: $m$ covers all tokens, optimizing both the revised reasoning chain and the safe response, allowing the model to learn the complete safe reasoning trajectory.
-Safety Backtrack: $m$ covers only the appended self-check and final response, while the original chain is masked as context, preserving the model’s reasoning distribution and enhancing self-reflection and safety.
-3.3 Workflow of COG
-Both SafR and SafB aim to enhance safety by mitigating certain aspects of Self-Jailbreak, following a three-phase structure.
-While their core workflows overlap, they differ in how the reasoning chain is modified and integrated.
-Phase 1: Input & Analysis (shared)
-Initial Inference. The base model $\pi_{0}$ infers on the harmful prompt $x$, producing the initial reasoning chain $c=[a(x),p(x)]$ and response $y$.
-Information Extraction. A structured prompt (see Appendix A.2) together with Qwen2.5-72B-Instruct (Team, 2024) extracts the risk analysis and response strategy.
-Failure Classification. Given $x$, $a(x)$, and $p(x)$, Qwen2.5-72B-Instruct determines whether the reasoning violates safety principles and assigns a Self-Jailbreak type.
-Phase 2: Data Construction (divergent)
-SafR – Reasoning Chain Recomposition and Merging.
-For each identified failure type, a tailored prompt guides $\pi_{0}$ to perform reasoning chain recomposition, adjusting the risk analysis and response strategy with safety-oriented considerations to generate corrected sub-chains.
-Subsequently, these recomposed analyses $a(x)$ and predictions $p(x)$ are merged into a coherent, Safety-oriented Chain-of-Thought (S-COT).
-SafB – Safety Backtrack Augmentation.
-Rather than recomposing, $\pi_{0}$ performs a reflective process on the original reasoning chain to generate a supplementary safety check $s(x)$.
-Carefully engineered transitional expressions ensure a natural integration between the reasoning chain and the check, resulting in a coherent S-COT.
-Phase 3: Inference & Alignment (shared, with minor variation)
-Safe Response Generation. Both SafR and SafB use the S-COT as the reasoning trajectory to guide $\pi_{0}$ in generating a safe response $y_{\text{safe}}$.
-Alignment Training. In both methods, $\pi_{0}$ is further trained on the safe response data, yielding the aligned and safety-enhanced model $\pi_{1}$ with strong reasoning capabilities.
-4 Experiment
-4.1 Experiment Setting
-Baselines We include four representative methods: STAR-1  (Wang et al., 2025b) and SafeChain  (Jiang et al., 2025) improve safety by performing SFT on a curated set of safe data.
-SAFEPATH  (Jeung et al., 2025) guides model behavior by training it to generate a fixed prefix, and SafeKey  (Zhou et al., 2025b) enforces stricter safety alignment by introducing two additional safety classifiers.
-Each method was reproduced using the officially released training configurations and datasets, with details provided in Appendix A.1.
-Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025).
-Our Safety Recomposition and Safety Backtrack methods are implemented via supervised fine-tuning (SFT) on the Qwen3 base models.
-As illustrated in Figure 3, the training and experimental procedures comprise three distinct phases.
-The details are shown in the Appendix A.2.2.
-Safety Benchmarks We evaluate model safety using four benchmarks.
-SORRY-Bench  (Xie et al., 2024) and StrongREJECT  (Souly et al., 2024) measure the model’s ability to refuse harmful prompts, while WildJailBreak  (Jiang et al., 2024b) and JailBreakBench  (Chao et al., 2024) assess its robustness against adversarial jailbreak attacks.
-[Uncaptioned image] Figure 4: Safety vs. reasoning trade-off for the 32B model.
-Reasoning Benchmarks To assess general reasoning ability, we use AIME2024  (Mathematical Association of America (2024), MAA) for mathematical reasoning and GPQA-Diamond  (Rein et al., 2024) for complex knowledge-based reasoning.
-To reduce randomness in evaluation, we perform multiple rollout rounds: GPQA is averaged over two rounds, and AIME over eight rounds.
-Additional details of the reasoning and safety benchmarks can be found in Appendix A.3.
-Training Dataset We extracted 15,000 high-quality harmful prompts from 7 public collections—Alert  (Tedeschi et al., 2024), ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming  (Ahmadian et al., 2024), Do-Not-Answer  (Wang et al., 2023), AttaQ  (Kour et al., 2023), and Toxic-Chat  (Lin et al., 2023).
-We then designated this curated set as our seed dataset and fed it into our pipeline to generate corresponding high-quality safety response sets.
-4.2 Main Result
-Table 2: Performance comparison of different methods under the main experimental setting.
-Here, lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.
+Safety Backtrack (SafB) SafB keeps the original reasoning chain and adds a targeted self-check step.
+Guided by the Self-Jailbreak type, the original model $\pi_{0}$ takes the original $a(x)$ and $p(x)$ and generates a self-check segment that focuses on the failure-inducing parts of the reasoning.
+This self-check revisits the earlier reasoning decisions and provides corrective guidance before producing the final response.
+We then append the self-check segment to the end of the original chain, forming an augmented, safety-oriented reasoning trajectory.
+3.4 Phase 3: Inference & Alignment
+In Phase 3, the safety-oriented COT produced by SafR or SafB is used to generate a final safe response, which is then used as training data for alignment.
+We then fine-tune the original model $\pi_{0}$ on this constructed data using selective loss masking.
+For SafR, we supervise the entire recomposed reasoning chain together with the final response.
+For SafB, we supervise only the appended self-check segment and the final response, while masking the original reasoning tokens.
+This selective supervision encourages learning safety-oriented corrections while mitigating unintended distributional shift away from $\pi_{0}$’s original reasoning trajectory.
 Method Harmful Jailbreak Reasoning
@@
 […]
-Vanilla 45.45 13.62 38.80 81.71 57.33 77.50
-STAR-1 18.86 0.74 80.00 62.2 57.33 71.25
-SAFEPATH 36.14 10.03 22.80 42.68 55.56 67.92
-SafeChain 49.55 16.99 36.80 70.73 52.28 66.25
-SafeKey 3.18 0.32 8.53 13.2 41.90 70.80
-Safety Backtrack(Ours) 17.05 2.05 8.00 26.83 54.30 77.50
-Safety Recomposition(Ours) 13.41 2.86 9.20 28.05 56.82 76.25
+Vanilla 45.45 13.62 38.80 81.71 29.00 57.33 77.50 97.6 92.68
+STAR-1 18.86 0.74 20.00 37.8 5.00 57.33 71.25 96.4 93.29
+SafePath 36.14 10.03 22.80 42.68 14.00 55.56 67.92 94.8 92.07
+SafeChain 49.55 16.99 36.80 70.73 39.00 52.28 66.25 95.2 90.24
+SafeKey 3.18 0.32 8.53 13.2 7.00 41.90 70.58 90.0 91.46
+Safety Backtrack (Ours) 17.05 1.45 8.00 26.83 5.00 54.30 77.50 97.4 93.90
+Safety Recomposition (Ours) 13.18 1.89 9.20 28.05 5.00 56.82 76.25 92.6 93.90
 […]
-Vanilla 55.45 12.44 34.00 68.29 63.14 77.92
-STAR-1 17.95 0.72 86.80 76.83 56.32 76.25
-SAFEPATH 34.09 8.49 16.00 50.00 57.78 70.42
-SafeChain 49.32 16.87 35.60 67.07 57.58 71.25
-SafeKey 4.77 0.32 4.88 10.40 49.00 76.70
-Safety Backtrack(Ours) 11.82 1.30 6.40 18.05 62.12 77.92
-Safety Recomposition(Ours) 12.50 4.97 2.80 21.95 60.36 78.75
+Vanilla 55.45 12.44 34.00 68.29 21.00 63.14 77.92 97.6 95.73
+STAR-1 17.95 0.72 13.20 23.17 32.00 56.32 76.25 88.4 97.56
+SafePath 24.09 8.49 16.00 50.00 15.00 57.78 70.42 78.8 88.41
+SafeChain 49.09 16.87 35.60 67.07 32.00 57.58 71.25 87.4 95.73
+SafeKey 4.77 0.32 4.88 10.40 6.00 49.00 76.70 88.4 86.58
+Safety Backtrack (Ours) 10.91 0.97 6.40 18.05 7.00 62.12 77.92 97.0 96.34
+Safety Recomposition (Ours) 7.95 2.09 2.80 21.95 5.00 60.36 78.75 97.4 96.34
 […]
-Vanilla 46.59 12.25 35.20 20.43 65.66 81.67
-STAR-1 18.41 0.74 16.80 35.37 54.55 72.92
-SAFEPATH 40.00 6.57 22.80 53.66 62.38 70.25
-SafeChain 47.95 16.39 28.40 70.73 54.30 71.70
-SafeKey 3.41 0.32 7.32 10.40 54.30 71.70
-Safety Backtrack(Ours) 14.55 2.77 8.80 23.17 61.62 77.08
-Safety Recomposition(Ours) 7.05 1.49 3.20 17.07 62.38 82.08
-Our results, summarized in Table 4.2, reveal several key findings:
-Baseline methods enhance safety performance at the substantial cost of reasoning capability.
-As reported in Table 4.2, several baselines either underperform the vanilla model on safety (e.g., SafeChain, STAR-1) or secure safety gains only at the cost of substantial reasoning degradation (e.g., SafeKey).
-COG achieves the best trade-off between safety and reasoning across model scales.
-Our proposed COG framework, including Safety Backtrack and Safety Recomposition, consistently achieves the strongest safety improvements while largely preserving reasoning ability.
-This trade-off is clearly visualized in Figure 4, where our methods are positioned in the optimal upper-right quadrant, demonstrating both high safety and high reasoning performance simultaneously.
-Specifically, across model scales, SafB and SafR attain the lowest or near-lowest harmful/jailbreak scores while maintaining competitive reasoning performance (GPQA-Diamond and AIME2024), highlighting the robustness and scalability of COG in resolving the safety–reasoning trade-off.
-5 Analysis
-5.1 PCA Analysis of Safety Improvements with Minimal Reasoning Impact
-To quantitatively evaluate the trade-offs between safety and reasoning, we apply Principal Component Analysis (PCA) to the hidden states of Qwen3-32B We analyze the geometric structure of the representation space under different training approaches, focusing on safety-critical prompts (Harmful, RedTeaming) and reasoning prompts (Reasoning).
-A linear decision boundary is trained on the two-dimensional PCA projection to quantify the separation between safety and reasoning  [^1] Results for the 8B and 14B models are provided in Appendix B.2., which is illustrated in Figure 5, where we can conclude that:
-Our approach establishes a more separable safety manifold while retaining the stability of reasoning representations, indicating effective safety gains with negligible reasoning degradation.
-Specifically, SafR achieves the largest safety improvement ($\Delta$_Harmful = +4.985) with minimal reasoning loss ($\Delta$_Reasoning = –0.676), whereas SafB provides a more balanced gain ($\Delta$_Harmful = +3.961, $\Delta$_Reasoning = –1.771).
-In contrast, SafeChain offers only marginal safety benefits ($\Delta$_Harmful = +0.378) while causing severe reasoning degradation ($\Delta$_Reasoning = –5.260).
-These results indicate that our COG framework substantially enhances safety while largely retaining reasoning capabilities, effectively improving safety alignment without disrupting core reasoning structures.
-Figure 5: PCA of the Qwen3-32B representation space.
-Gray ellipses indicate base model distributions, while colored regions denote post-training clusters (red: safety, blue: reasoning).
-Centroid shifts ($\Delta$) quantify displacement relative to the linear decision boundary (dashed), with $\Delta>0$ indicating increased margin and $\Delta<0$ indicating decreased margin.
-5.2 Understanding Reasoning Preservation via Trajectory and Token Analysis
-To study how training strategies influence the reasoning structure of language models, we analyze their cognitive trajectories using DeepSeek-v3.1 (DeepSeek-AI, 2024) with established prompts (Zeng et al., 2025; Gandhi et al., 2025), and measure how these patterns differ from a base model.
-We track the frequency of key reasoning behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—with scores reflecting the average occurrences per problem on the AIME benchmark, which is presented in Table 3, which indicates that:
-Models trained with our methods best maintain the cognitive structure of the base model, and therefore have only a minimal impact on the model’s reasoning capabilities.
-Specifically, SafR exhibits slightly higher behavior frequencies (+0.3%) and SafB slightly lower (–0.3%), indicating that both methods achieve high alignment with the base model’s reasoning patterns.
-In contrast, alternative methods such as SafeChain, SafePath, and star-1 exhibit larger deviations (see Table 3).
-Furthermore, the token-level analysis in Appendix B.3 confirms that SafR and SafB maintain token usage close to the Base model while improving safety, indicating that they enhance safety without significantly altering intrinsic reasoning behavior.
-Table 3: Comparison of reasoning patterns (Qwen3-32B) across different training strategies.“Overall Avg.” denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
-Pattern Base SafeChain SafePath star-1 SafB SafR
+Vanilla 46.59 12.25 35.20 80.43 23.00 65.66 81.67 97.6 98.17
+STAR-1 18.41 0.83 16.80 35.37 3.00 54.55 72.92 85.2 95.12
+SafePath 40.00 6.57 22.80 53.66 20.00 62.38 70.25 60.4 97.56
+SafeChain 47.95 16.39 28.40 70.73 34.00 54.30 71.70 86.4 96.34
+SafeKey 3.41 0.32 7.32 10.40 10.00 54.30 71.70 86.8 87.20
+Safety Backtrack (Ours) 14.55 1.68 8.80 23.17 1.00 61.62 77.08 97.4 98.17
+Safety Recomposition (Ours) 7.05 1.10 3.20 17.07 2.00 62.38 82.08 97.6 97.56
+Table 2: Performance comparison of different methods under the main experimental setting.
+Lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.
+4 Experiment
+In this section, we conduct experiments across multiple safety and reasoning benchmarks to validate the effectiveness of our proposed COG framework.
+We demonstrate that compared to prior baselines, our methods achieve state-of-the-art safety-reasoning balance across different model scales.
+Furthermore, we perform in-depth analyses to investigate the underlying mechanisms that contribute to the success of our approach, examining both the preservation of reasoning patterns and the distributional characteristics of the learned representations.
+4.1 Experiment Setting
+Training Dataset We collect 15,000 high-quality harmful queries from public datasets including Alert, ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming, Do-Not-Answer, AttaQ, and Toxic-Chat (Tedeschi et al., 2024; Ahmadian et al., 2024; Wang et al., 2023; Kour et al., 2023; Lin et al., 2023).
+This curated set serves as input to our pipeline for constructing corresponding safety-aligned responses.
+Evaluation Benchmarks For safety, we use Sorry-bench and StrongREJECT for harmful-prompt refusal, and WildJailBreak and JailBreakBench for jailbreak robustness (Xie et al., 2024; Souly et al., 2024; Jiang et al., 2024b; Chao et al., 2024).
+For reasoning, we adopt GPQA-Diamond, AIME2024, MATH500, and HumanEval to assess math and code reasoning capabilities (Rein et al., 2024; Mathematical Association of America (MAA), 2024; Lightman et al., 2023; Chen et al., 2021).
+Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025). SafR and SafB are implemented with training details in Appendix A.2.1.
+Baselines We compare with diverse representative safety alignment baselines including STAR-1, SafeChain, SafePath, and SafeKey (Wang et al., 2025; Jiang et al., 2025; Jeung et al., 2025; Zhou et al., 2025b).
+All baselines are reproduced following official codes, configurations, and datasets, detailed in Appendix A.2.1
+4.2 Overall Results
+Table 3.4 summarizes the overall safety and reasoning performance across model scales.
+The results reveal clear differences in how existing methods trade off safety and reasoning, and highlight the effectiveness of CoG under this trade-off.
+Figure 5: Safety vs. reasoning trade-off for the Qwen 32B.
+1) Vanilla LRMs without safety alignment exhibit poor safety performance.
+Despite their impressive performance on reasoning benchmarks, the vanilla versions of Qwen3-8B, 14B, and 32B consistently achieve the highest (worst) scores across all safety and jailbreak metrics.
+For instance, on the Sorry-bench and Wildjailbreak benchmarks, the vanilla models demonstrate a high propensity to follow harmful instructions, with Sorry-bench scores reaching as high as 55.45% in the 14B variant.
+This confirms that advanced reasoning capabilities do not inherently translate to safety, and without explicit alignment, these models remain fragile.
+2) Previous approaches enhance safety performance at the substantial cost of reasoning capability.
+Baselines such as SafeKey and STAR-1 effectively reduce safety risks, but often incur a clear compromise in reasoning.
+For instance, on Qwen3-32B, SafeKey reduces Sorry-bench from 46.59% to 3.41% and JBB-PAIR from 80.43% to 10.40%, at the expense of lowering GPQA-Diamond from 65.66% to 54.30% and AIME2024 from 81.67% to 71.70% Similar trade-offs are observed for STAR-1 across model scales, suggesting that the global safety signal is harmful for reasoning capability.
+3) COG achieves the best safety-reasoning balance across all model scales.
+As shown in Table 3.4, our methods significantly boost safety, reducing Wildjailbreak and Sorry-bench scores to levels comparable with the strongest safety baselines.
+This trade-off is clearly visualized in Figure 5, where our methods achieve the most favorable balance between safety and reasoning performance.
+For example, on Qwen3-32B, Safety Recomposition reduces Sorry-bench from 46.59% to 7.05% while maintaining AIME2024 at 82.08% and MATH500 at 97.6% compared to the vanilla model; Safety Backtrack provides a complementary point that further preserves reasoning with slightly weaker safety than the most safety baseline.
+Overall, these results indicate that CoG can deliver strong safety gains with minimal loss in reasoning ability, enabling a more favorable safety-reasoning trade-off for deploying capable LRMs.
+4.3 Detailed Analysis
+Pattern Vanilla SafeChain SafePath Star-1 SafB SafR
 Backtracking 1.33 1.10 1.20 1.30 1.27 1.30
@@
 […]
-Table 4: Comparison of model performance with and without masking the default reasoning trajectory.
-“SafB (full mask)” corresponds to the Safety Backtrack (full mask) setting, while “SafB (part.
-mask)” corresponds to the Safety Backtrack (partial mask) setting.
+Table 3: Comparison on the frequencies of reasoning patterns (Qwen3-32B) across different training strategies.
+“Overall Avg." denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
+Figure 6: PCA of the Qwen3-32B representation space.
+Gray ellipses indicate base model distributions, while colored regions denote post-training clusters (red: safety, blue: reasoning).
+Centroid shifts ($\Delta$) quantify displacement relative to the linear decision boundary (dashed), with $\Delta>0$ indicating increased margin and $\Delta<0$ indicating decreased margin.
+We further examine why COG improves safety without sacrificing reasoning largely.
+Specifically, we test whether COG preserves the base model’s reasoning paradigm from two perspectives—reasoning trajectory patterns and representation separability —and provide illustrative case studies in Appendix D.
+4.3.1 Reasoning Pattern
+To study how training strategies influence the reasoning structure of LRMs, we begin with a simple token-level signal.
+Based on our token-level statistics, SafR and SafB induce minimal changes in generation length on GPQA-Diamond and AIME compared to the base model, while baselines—such as STAR-1—consistently produce substantially shorter outputs across all model scales; see Appendix C.2 for details.
+Subsequently, we conduct a deeper reasoning-pattern analysis, which calculates the frequencies of reasoning patterns among various methods.
+Following established prompting protocols (Zeng et al., 2025; Gandhi et al., 2025), we use DeepSeek-v3.1 (DeepSeek-AI, 2024) to analyze cognitive trajectories and quantify four key behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—measured as the average occurrences per problem on AIME.
+CoG resembles the most to original LRMs in reasoning frequencies.
+Table 3 indicates that SafR and SafB remain the closest to the vanilla model in behavior frequencies (SafR: +0.03%, SafB: -0.03%), while SafeChain, SafePath, and STAR-1 exhibit substantially larger deviations.
+Overall, these results suggest that CoG improves safety while largely preserving original reasoning patterns.
+4.3.2 Representation Separability
+To quantify safety-reasoning separability at the representation level, we apply PCA to Qwen3-32B hidden states on safety-critical prompts and reasoning prompts, and fit a linear decision boundary to quantify their separability, as illustrated in Figure 6.
+We define each cluster’s margin as the distance from its centroid to the decision boundary: a larger margin means the centroid is farther from the boundary and thus easier to separate, while a smaller margin indicates weaker separability.
+CoG improves safety separation while largely preserving reasoning representations.
+SafR produces the largest increase in the safety margin (+4.985+4.985) with only a small reduction in the reasoning margin (0.676-0.676).
+SafB increases the safety margin (+3.961+3.961), but reduces the reasoning margin more (1.771-1.771), placing reasoning representations closer to the decision boundary.
+In contrast, SafeChain yields a small safety gain (+0.378+0.378) while greatly reducing the reasoning margin (5.260-5.260), indicating much weaker separability for reasoning representations.
+Overall, these results suggest CoG strengthens safety separation without strongly distorting reasoning representations.
+Comparison Awareness Classification
+Human A 0.81 0.77
+Human B 0.79 0.71
+Human C 0.78 0.73
+Mean Consensus 0.79 0.74
+Table 4: Human–model agreement (Pearson’s $r$) on 50 instance-level binary judgments for Self-Jailbreak classification and LlamaGuard, where Awareness stands for the awareness of the risk in queries, while Classification represents the classification of Self-Jailbreak.
+4.3.3 Consistency with human
+To verify the effectiveness of LLM-as-a-judge in CoG framework and discovering Self-Jailbreak, we conduct a human–model agreement on a sampled subset from Wildjailbreak by comparing human judgments with model decisions when deciding the awareness of the risk in query as well as the classification of Self-jailbreak.
+As shown in Table 4, both systems exhibit strong and consistent alignment with human judgments across annotators, achieving correlations of 0.74 (Awareness) and 0.79 (Classification) with the averaged human consensus.
+4.4 Ablation Study
+To verify the effect of the selective loss masking strategy in SafB, we compare (i) partial-mask training (our default), where supervision is applied only to the self-check segment and the final answer, and (ii) no-mask training, where the full sequence (including the original reasoning trace) is supervised.
 […]
-Qwen3-8B 45.45 13.62 38.80 81.71 57.58 73.30
-SafB (full mask) 17.05 2.05 8.00 26.83 54.30 77.50
-SafB (part. mask) 23.64 5.37 22.40 59.76 56.87 75.00
-5.3 Impact of Masking Strategy in Safety Backtrack
-We evaluate the masking strategy in Safety Backtrack (SafB) by comparing two training setups: (1) Full-mask training, where the loss is applied to the reasoning trace, self-check, and final answer; and (2) Partial-mask training (our main method), where the loss is applied only to the self-check and final answer, with the reasoning trace masked.
-Table 4 shows that both SafB variants improve safety, with partial masking yielding larger gains on the Harmful and RedTeaming benchmarks.
-For reasoning, both variants enhance AIME, whereas GPQA shows a slight degradation, more pronounced under full masked SafB, which can be attributed to fewer rollout samples (two for GPQA vs.
-eight for AIME).
+SafB 17.05 2.05 8.00 26.83
+SafB (w/o mask) 23.64 5.37 22.40 59.76
+Table 5: Effect of selective loss masking in SafB.
+Table 5 shows that selective masking is crucial for SafB.
+Partial-mask training consistently achieves better results on safety benchmarks.
+In contrast, removing the mask and supervising the full reasoning trace substantially degrades safety, with the largest drops observed on jailbreak benchmarks.
+This supports masking the original reasoning trace in SafB and applying supervision only to the self-check and final answer, which helps mitigate unintended distributional shift.
 5 Related Work
-Safety–Reasoning Trade‐Off
-Reasoning capability and safety appear to be inherently conflicting objectives in LRMs.
-Zhou et al.
-(Zhou et al., 2025a) show that stronger LRMs tend to produce more unsafe content, while studies by Huang et al.
-(Huang et al., 2025) and Li et al.
-(Li et al., 2025) highlight a fundamental trade‐off: improving one objective inevitably undermines the other, underscoring the need to balance safety and reasoning performance.
-Safety for Reasoning
-Instead of relying on external filters, safety-for-reasoning methods integrate controls directly into inference through training.
-Current approaches mostly build on supervised fine-tuning (SFT).
-STAR-1  (Wang et al., 2025b), SafeChain  (Jiang et al., 2025), SAFEPATH  (Jeung et al., 2025), and SafeKey  (Zhou et al., 2025b) apply curated data and tailored SFT to enhance safety while retaining reasoning ability.
-These methods better preserve coherence but still deliver only modest safety gains or impose notable performance costs, leaving room for more efficient optimization.
+Vulnerable Safety of LRMs.
+Recent studies consistently show that LRMs remain vulnerable under harmful queries and adversarial settings, and that longer, explicit reasoning can introduce extra safety risks.
+(Huang et al., 2025) shows a clear safety-reasoning trade-off in common alignment pipelines, where better safety can hurt reasoning performance.
+(Zhu et al., 2025) shows that attackers can manipulate special delimiter markers to bypass reasoning, effectively skipping the intended deliberation.
+(Xu et al., 2025) demonstrates that safety alignment can be undermined via fine-tuning attacks against CoT-enabled models.
+Other work, such as (Li et al., 2025; Zhang et al., 2025a), also studies how safety changes as reasoning improves, and shows that reasoning-time safety can still fail in hard or adversarial cases.
+Safety Alignment for LRMs.
+To mitigate these risks, recent methods align LRMs by shaping the reasoning process or injecting safety deliberation.
+(Guan et al., 2024) trains models to explicitly recall and reason over safety specifications before answering.
+Data-driven approaches such as (Wang et al., 2025; Zhang et al., 2025a) construct safety-oriented reasoning trajectories for fine-tuning.
+(Jiang et al., 2025) both assesses long-CoT safety risks and introduces CoT-style safety training data.
+Orthogonally, (Zhou et al., 2025b) aims to strengthen internal safety activation signals, while (Doula et al., 2025) promotes early safety priming to reduce harmful reasoning.
+In contrast, we analyze reasoning trajectories stage by stage and formalize Self-Jailbreak as a failure mode after risk awareness.
 6 Conclusion
-In this paper, we conducted a systematic investigation into the safety mechanisms of LRMs.
-Our analysis reveals the existence of Self-Jailbreak, where LRMs persuade themselves into answering the harmful queries.
-This phenomenon sheds new light on the underlying mechanisms of the reasoning process in LRMs.
-Building on this discovery, we developed CoG, a training framework that suppresses Self-Jailbreak by detecting and mitigating unsafe reasoning patterns through Safety Backtrack and Safety Recomposition.
-Experimental results show that CoG substantially enhances safety with minimal degradation in reasoning capability, achieving state-of-the-art performance in the safety–reasoning trade-off.
-Overall, this study provides a principled framework for aligning LRMs toward safe and reliable reasoning.
+We systematically analyze safety failures in LRMs and uncover Self-Jailbreak, in which the model initially recognizes harmful intent but later overrides this judgment during reasoning, leading to unsafe outputs.
+This finding suggests that many safety failures stem from failure-inducing steps that override correct risk awareness within the reasoning chain.
+Motivated by this, we propose Chain-of-Guardrail (CoG), a trajectory-level training framework that mitigates Self-Jailbreak via targeted, step-level interventions.
+Across multiple safety and reasoning benchmarks, CoG improves safety and robustness while maintaining comparable reasoning performance.
+Overall, our study provides an analysis lens and a practical framework for aligning LRMs toward safer reasoning.
+Limitations
+We acknowledge several limitations in the current study.
+First, due to computational resource constraints, we do not evaluate our method on substantially larger-scale reasoning models.
+While our experiments demonstrate the effectiveness of the proposed framework on the models studied in this paper, validating its behavior at larger scales remains an important direction for future work.
+Second, a large portion of our evaluation relies on automated judgments using LLM-based evaluators.
+Although LLM-as-judge provides scalability and consistency for analyzing fine-grained reasoning behaviors, it remains an imperfect proxy for human judgment.
+To partially mitigate this limitation, we conduct human–model consistency analyses and observe strong agreement between human annotations and automated evaluations.
+Nevertheless, establishing more reliable, standardized, and cost-effective evaluation protocols for safety failures in long-form reasoning remains an open problem.
+We leave the development of improved human-in-the-loop or hybrid evaluation frameworks to future work.
 Appendix A Detailed Experimental Setup
@@
 SafePath
-SAFEPATH fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first” at the start of inference, guiding the LRMs to consider more about safety during the generation process.
+SafePath fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first" at the start of inference, guiding the LRMs to consider more about safety during the generation process.
 SafeKey
@@
 We use the official dataset and replicate the experiments following the parameter settings reported in the original paper. The detailed training configurations are presented in Table 6.
-Table 5: Detailed Training Hyperparameters for Star-1
-Hyperparameter Value Hyperparameter Value
-Finetuning Type Full Optimizer AdamW
-Adam $\beta_{1},\beta_{2}$ 0.9, 0.95 Learning Rate 1e-5
-Epochs 5.0 Batch Size 2
-Gradient Accumulation Steps 8 Weight Decay 1e-4
-Warmup Ratio 0.05 Cutoff Length 8,192
+Hyperparameter Value
+Finetuning Type Full
+Optimizer AdamW
+Adam $\beta_{1},\beta_{2}$ 0.9, 0.95
+Learning Rate 1e-5
+Epochs 5.0
+Batch Size 2
+Gradient Accumulation Steps 8
+Weight Decay 1e-4
+Warmup Ratio 0.05
+Cutoff Length 8,192
+Table 6: Detailed training hyperparameters for Star-1.
 SafeChain
 We trained Qwen3 series models with the original SafeChain dataset with llama-factory. Detailed implementation of SafeChain experiment is described as shown in Table 7:
-Table 6: Detailed Training Hyperparameters for SafeChain
-Parameter Value Parameter Value
-Epochs 2 Batch Size 2
+Parameter Value
+Epochs 2
+Batch Size 2
 Gradient Accumulation Steps 2
+Table 7: Detailed training hyperparameters for SafeChain.
 Safekey
 We use the official SafeKey codebase, making only model-level modifications to its startup scripts. The detailed implementation of the SafeKey experiment is described as shown in Table 8.
-Table 7: Detailed Training Hyperparameters for SafeKey
-Parameter Value Parameter Value
-Epochs 5 Batch Size 2
+Parameter Value
+Epochs 5
+Batch Size 2
 Gradient Accumulation Steps 8
+Table 8: Detailed training hyperparameters for SafeKey.
 SafePath
-The detailed implementation of the SafePath experiment is described in Table 8:
-Table 8: Detailed Training Hyperparameters for SafePath
-Parameter Value Parameter Value
-Finetuning Type Full Cutoff Length 8192
-Batch Size 2 Gradient Accumulation Steps 2
-Learning Rate 1e-5 Max Steps 20
+The detailed implementation of the SafePath experiment is described in Table 9.
+Parameter Value
+Finetuning Type Full
+Cutoff Length 8192
+Batch Size 2
+Gradient Accumulation Steps 2
+Learning Rate 1e-5
+Max Steps 20
 Warmup Ratio 0.05
+Table 9: Detailed training hyperparameters for SafePath.
 A.2 Implementation Details of Our Method
-A.2.1 Manual Classification Accuracy
-Table 9: Classification Accuracy of Three Graduate Students on 50 Cases
-Participant Correct Cases Accuracy (%)
-Graduate Student A 46 92.0
-Graduate Student B 43 86.0
-Graduate Student C 44 88.0
-Average 44.33 88.67
-We invited three graduate students with domain expertise to classify 50 representative cases manually.
-As shown in Table 9, their classification accuracy ranged from 86.0% to 92.0%, with an average of 88.67%, indicating the reliability of our labeling scheme.
-A.2.2 COG Generation Parameters
+A.2.1 COG Generation Parameters
 During the sampling process(Phase 1), to ensure output diversity and prevent model degeneration, we set the temperature to 0.7, top_p to 0.8, and presence_penalty to 1.5 to produce the original responses used as seed data (see Table 10).
@@
 These carefully chosen parameters balance generation quality and diversity while minimally impacting the model’s reasoning capability.
-Table 10: Generation Parameter Settings
-Stage Parameters Values
+Stage Parameter Value
 Generation Phase temperature 0.7
@@
 presence_penalty 1.5
-Table 11: Training Hyperparameters
+Table 10: Generation parameter settings.
+Parameter Value
+Finetuning Type Full
+Learning Rate 2e-6
+Cutoff Length 8192
+Epochs 3.0
+Batch Size 2
+Warmup Ratio 0.1
+Gradient Accumulation Steps 4
+Table 11: Training hyperparameters.
+Safety Benchmarks Reasoning Benchmarks
 Parameter Value Parameter Value
-Finetuning Type full Learning Rate 2e-6
-Cutoff Length 8192 Epochs 3.0
-Batch Size 2 Warmup Ratio 0.1
-Gradient Accumulation Steps 4
-Table 12: Benchmark Implementation Details
-Safety Benchmarks Reasoning Benchmarks
-Parameters Values Parameters Values
 temperature 0.7 temperature 0.6
-top_p 1 top_k 20
+top_p 1.0 top_k 20
 max_new_tokens 16384 top_p 0.95
@@
 AIME2024 rollout 8
+MATH500 rollout 1
+HumanEval rollout 1
+Table 12: Benchmark implementation details.
 COG Training Parameters
@@
 The detailed hyperparameter setting is shown in Table 12.
-Appendix B Additional Experimental Results
-B.1 Initial Investigation by Switching Thinking Mode
-Recent works  (Zhou et al., 2025a; Zhang et al., 2025) have discovered that LRMs tend to answer harmful questions.
-To further validate whether the thinking trajectory influences the safety of LLMs, we compare multiple LRMs with the thinking mode switch on & off.
-The evaluation is performed on 2k harmful queries from Wildjailbreak  (Jiang et al., 2024a), with safety assessed using Llama-Guard-3-8B (Llama Team, 2024) as the automatic judge.
-The thinking trajectory substantially undermines the safety of LRMs.
-As shown in Table 13, LRMs become consistently safer when the thinking mode is disabled, but their safety degrades notably once it is enabled.
-This trend holds across different model scales: enabling the thinking trajectory systematically increases the likelihood of harmful responses.
-For example, the harmful answer rate of Qwen3-32B rises by 10.55% when the mode is enabled, while the 8B and 14B variants also exhibit noticeable increases (6.04% and 3.35%, respectively).
-These results indicate that content generated within the thinking trajectory can directly drive LRMs to produce harmful outputs.
-Table 13: Percentage of harmful answers (”Answer”) and refusals (”Refuse”) on Wildjailbreak. Higher ’Answer’ indicates lower safety. $\Delta=$ Disabled - Enabled.
-Model Enabled Disabled $\Delta$ (Answer)
-Answer Refuse Answer Refuse
-Qwen3-8B 41.40 58.60 35.36 64.64 6.04-6.04
-Qwen3-14B 35.65 64.35 32.30 67.70 3.35-3.35
-Qwen3-32B 41.45 58.55 30.90 79.10 10.55-10.55
-B.2 Analysis: PCA Analysis of 8B and 14B Models
+Appendix B Initial Investigation by Switching Thinking Mode
+Recent works (Zhou et al., 2025a; Zhang et al., 2025b) have observed that LRMs tend to answer harmful questions.
+To further examine whether the thinking trajectory causally influences model safety, we evaluate multiple LRMs with the thinking mode switched on and off.
+Specifically, we test 2k harmful queries from WildJailbreak (Jiang et al., 2024a), and use Llama-Guard-3-8B (Llama Team, 2024) as the automatic safety judge.
+Figure 7 reports the proportion of harmful answers under the two settings across Qwen3 variants.
+A consistent pattern emerges across model scales: enabling the thinking mode substantially increases the likelihood of producing harmful answers.
+For all three Qwen3 variants, turning on thinking leads to a higher harmful-answer rate, with increases from 35.4% to 41.4% (8B), 32.3% to 35.6% (14B), and 30.9% to 41.5% (32B).
+Notably, this effect does not diminish with model scaling.
+In fact, the largest model exhibits the most pronounced gap, suggesting that stronger reasoning capacity alone does not guarantee safer behavior.
+Instead, explicit reasoning may create additional opportunities for models to rationalize or justify unsafe responses.
+Collectively, these results offer preliminary evidence that explicit reasoning trajectories may amplify safety risks, motivating a finer-grained investigation of how unsafe behavior is instantiated within intermediate reasoning steps rather than being determined solely by the final output.
+Figure 7: Harmful answer rate (%) of Qwen3 models on 2k harmful queries from WildJailbreak, comparing thinking mode enabled vs.
+disabled.
+Safety is automatically judged by Llama-Guard-3-8B.
+Enabling thinking consistently increases the proportion of harmful answers across model scales.
+Appendix C Additional Experimental Results
+C.1 Analysis: PCA Analysis of 8B and 14B Models
 We conducted a comparative analysis of models with different parameter sizes and fine-tuning methods (SafR and SafB), aiming to evaluate their impact on safety and representational clustering.
-The observed differences are visualized in Figure 6 and 7, while quantitative results across all configurations are reported in Table 14.
-Table 14: Safety distance analysis of Qwen3 models by model size.
-Values represent perpendicular distances from cluster centroids to the linear decision boundary.
-Positive $\Delta$ values indicate movement away from the boundary (greater margin, safer region), while negative $\Delta$ values indicate movement toward the boundary (reduced margin, less safe region).
-Method Qwen3-8B Qwen3-14B Qwen3-32B
-Harmful Reasoning Harmful Reasoning Harmful Reasoning
-$\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$ $\Delta$
-Base 13.24 14.48 15.54 9.05 11.20 12.18
-SafR 18.55 15.02 18.95 8.33 16.18 11.50
-+5.30 +0.54 +3.41 -0.73 +4.98 -0.68
-SafB 19.51 12.95 19.14 9.63 15.16 10.41
-+6.27 -1.54 +3.60 +0.58 +3.96 -1.77
-SafeChain 13.45 5.87 12.59 4.26 11.58 6.92
-+0.21 -8.61 -2.95 -4.79 +0.38 -5.26
+The observed differences are visualized in Figure 8 and 9, while quantitative results across all configurations are reported in Table 7.
+Figure 8: PCA of the Qwen3-8B representation space.
+Figure 9: PCA of the Qwen3-14B representation space.
 Model Scale
@@
 Fine-Tuning Methods
-Both Safety Recomposition (SafR) and Safety Backtrack (SafB) fine-tuning strategies contribute to enhanced model safety, as evidenced by the centroid shifts observed in Figure 6 and Figure 7, and further quantified by the increase in Safety Distance (e.g., for the 8B model: from 20.05 to 25.68 after SafR).
-Notably, SafR yields consistently superior clustering performance, with higher Silhouette Scores than SafB across both model sizes.
-This suggests that SafR not only improves safety alignment but also leads to more coherent and well-separated internal representations.
-Figure 6: PCA of the Qwen3-8B representation space.
-Figure 7: PCA of the Qwen3-14B representation space.
-B.3 Preservation of Reasoning Token Usage
-Table 15 reports the average token counts for GPQA and AIME benchmarks across different training methods.
-SafR and SafB maintain token usage close to the Base model, while improving safety.
-In contrast, other baselines exhibit larger deviations, suggesting potential inefficiencies in reasoning.
-These results demonstrate that SafR and SafB enhance model safety without significantly affecting intrinsic reasoning behavior, selectively reinforcing safety-related features while preserving core reasoning capabilities.
-Table 15: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3-8B, Qwen3-14B, and Qwen3-32B across different training methods.
-Model Method GPQA-Diamond AIME
-Qwen3-8B Base 7553.84 14895.49
+Both Safety Recomposition (SafR) and Safety Backtrack (SafB) substantially improve model safety, as reflected by the centroid shifts in Figure 8 and Figure 9 and by the corresponding increases in Safety Distance reported in Table C.2.
+SafB reliably pushes the harmful cluster farther from the decision boundary while keeping the reasoning cluster relatively stable, indicating a balanced improvement in safety with limited impact on reasoning.
+C.2 Preservation of Reasoning Token Usage
+Table C.2 reports the average token length on GPQA-Diamond and AIME across training methods.
+Overall, SafR and SafB produce token counts comparable to the Base model, suggesting that their safety gains are not achieved by shortening generations.
+For example, on Qwen3-32B, the Base model averages 5150.81/12634.11 tokens (GPQA/AIME), while SafR remains close at 4830.90/12548.96 and SafB at 6452.00/12318.87.
+In contrast, SafePath shows a pronounced reduction to 3925.96/10200.49, with similar downward shifts on Qwen3-8B and Qwen3-14B.
+These results are consistent with the view that our methods improve safety while largely preserving the model’s intrinsic reasoning behavior.
+Method GPQA-Diamond AIME
+[…]
+Base 7553.84 14895.49
 SafePath 4869.04 11816.52
@@
 SafB 6540.14 14881.68
-Qwen3-14B Base 5585.28 14419.94
+[…]
+Base 5585.28 14419.94
 SafePath 4379.39 10939.46
@@
 SafB 5091.59 13010.36
-Qwen3-32B Base 5150.81 12634.11
+[…]
+Base 5150.81 12634.11
 SafeChain 5588.61 11198.09
@@
 SafR 4830.90 12548.96
-B.4 Analysis: Reasoning Safety Analysis
-Prior work has primarily focused on improving the safety of reasoning through external interventions (safety for reasoning).
-In this section, we take the opposite view and explore whether a model’s internal reasoning ability can be leveraged to enhance overall safety (reasoning for safety).
-To evaluate our approach, we compare it against the recently proposed R2D method  (Zhu et al., 2025) as a representative baseline.
-The experimental results are shown in Table 16.
-Table 16: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline. Lower is better for SafB, SafR, W-JB, and JB-B; higher is better for GPQA and AIME.
-Model SafB SafR W-JB JB-B GPQA AIME
+Table 13: Average token length on GPQA-Diamond and AIME benchmarks for Qwen3 models across different training methods.
+Method Harmful $\Delta$ Reasoning $\Delta$
+[…]
+Base 12.996 0.000 13.983 0.000
+SafB 19.615 +6.619 12.800 –1.183
+SafR 19.615 +6.619 12.800 –1.183
+SafeChain 13.609 +0.613 5.324 –8.659
+[…]
+Base 15.540 0.000 9.054 0.000
+SafB 19.142 +3.602 9.632 +0.578
+SafR 18.952 +3.412 8.328 –0.726
+SafeChain 12.590 –2.951 4.265 –4.789
+[…]
+Base 11.197 0.000 12.180 0.000
+SafB 15.158 +3.961 10.409 –1.771
+SafR 16.182 +4.985 11.504 –0.676
+SafeChain 11.576 +0.379 6.921 –5.259
+Table 14: Safety distance and relative changes ($\Delta$) for Qwen3 models under different safety training methods.
+Model S-B S-RB W-JB JB-B GPQA AIME
 Qwen3-8B 45.45 13.62 38.80 18.29 57.58 73.30
@@
 SafB(Ours) 17.05 2.05 8.00 73.17 54.30 77.50
-Based on the experimental results shown in the Table 16, we observe that applying reasoning-for-safety methods to LRMs indeed improves safety performance.
-However, similar to SafeKey and Star-1, these methods tend to sacrifice reasoning ability, failing to strike a proper balance between safety and reasoning capabilities.
-Although our methods partially address this issue, maintaining the integrity of reasoning ability remains a significant challenge.
-Appendix C Self-Jailbreak Category Example
-C.1 Benign Reframing
+Table 15: Performance comparison on safety and reasoning benchmarks between our COG and the R2D baseline.
+S-B denotes Sorry-bench, S-RB denotes StrongREJECT, W-JB denotes WildJailbreak, and JB-B denotes JailBreak Bench.
+Lower is better for S-B, S-RB, W-JB, and JB-B; higher is better for GPQA and AIME.
+C.3 Preservation of Reasoning Token Usage
+To further contextualize the safety–reasoning trade-off, we contrast two representative paradigms: (i) reasoning-for-safety (or reasoning-as-defense), which improves robustness by explicitly injecting safety-aware reasoning signals (e.g., reflection or step-level self-assessment), and (ii) safety-for-reasoning, which aims to enhance safety while preserving the integrity of an LRM’s original reasoning process.
+Table 15 compares our COG-based variants with a representative reasoning-as-defense baseline, R2D.
+While R2D achieves strong gains on safety-related metrics (SafB/SafR and WildJailbreak-style benchmarks), it exhibits a substantial drop in reasoning performance on GPQA and AIME, indicating that heavy safety-oriented reasoning control may interfere with general reasoning quality.
+In contrast, our SafB/SafR maintain markedly stronger reasoning performance (GPQA/AIME) while remaining competitive on safety benchmarks, suggesting that trajectory-level interventions can better preserve reasoning behavior under safety constraints.
+Nevertheless, the results also indicate that fully preserving reasoning integrity while maximizing safety remains challenging, motivating further study on how safety constraints reshape intermediate reasoning trajectories.
+Appendix D Case Study: COT Structure Stability
+To further demonstrate that our method maintains structural consistency with the original model’s reasoning patterns, we conduct a case study on a problem from AIME 2024.
+As shown in Figure 10, SafR exhibits a highly consistent reasoning structure with Qwen3-8B, while Star-1 employs a fundamentally different approach and ultimately produces an incorrect answer.
+Specifically, as highlighted in blue, both Qwen3-8B and SafR decompose the problem into two symmetric cases based on the starting direction, apply the same combinatorial formula $C(7,2)\times C(7,1)=147$ for each case, and correctly aggregate the results as $147+147=294$.
+In contrast, Star-1 (red highlights) introduces a "four types" classification strategy, leading to systematic double-counting and yielding $588=2\times 294$.
+This case illustrates that SafR preserves not only the base model’s reasoning steps but also its mathematical correctness.
+Appendix E Example of Safety Failure in LRMs
+E.1 Harm Misidentification
+The example of Warning in Harm Misidentification is presented in Figure 14, where LRM answers the harmful quesiton as the regular one.
+E.2 Self-Jailbreak: Benign Reframing
 The example of Benign Reframing in Self-Jailbreak is presented in Figure 11, where red text represents their Self-Jailbreak act.
-C.2 Logical Fallacies
+E.3 Self-Jailbreak: Logical Fallacies
 The example of Logical Fallacies in Self-Jailbreak is presented in Figure 12, where red text represents their Self-Jailbreak act.
-C.3 Warning
+E.4 Self-Jailbreak: Warning
 The example of Warning in Self-Jailbreak is presented in Figure 13, where red text represents their Self-Jailbreak act.
-C.4 Harm Identification
-The example of Warning in Harm Identification is presented in Figure 11, where LRM answers the harmful quesiton as the regular one.
-Appendix D Prompt Design
+Appendix F Prompt Design
 In this section, we present the prompt designs used in our COG framework and other experiments.
-D.1 Extraction Prompt
+F.1 Extraction Prompt
 This subsection provides both the base prompts and few-shot examples used in the extraction task.
@@
 We also include few-shot examples to guide the model during extraction, as illustrated in Figure 16.
-D.2 Classification Prompt
+F.2 Classification Prompt
 Prompt
 The prompt used for the classification task is outlined below in Figure 17.
-D.3 Safety Recomposition Prompt
+F.3 Safety Recomposition Prompt
 For the Safety Recomposition stage, our prompt is constructed by concatenating several components.
-Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts” field in a JSON file.
-Then, we combine the ”main_prompt”, the selected ”sub_prompt”, and a ”format_prompt” to form the final prompt.
+Specifically, based on the classification result, we extract a corresponding rewrite instruction from the “sub_prompts" field in a JSON file.
+Then, we combine the "main_prompt", the selected "sub_prompt", and a "format_prompt" to form the final prompt.
 The detailed structure is as shown in Figure 18
-D.4 Safety Backtrack Prompt
+F.4 Safety Backtrack Prompt
 In the Safety Backtrack stage, we further extend the prompt structure from the Safety Recomposition process.
-In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases”, selected based on the classification result.
+In addition to the previous components, we incorporate a transition phrase from “contextual_transition_phrases", selected based on the classification result.
 This helps guide the model more smoothly and maintain coherence in the final output.
 The detailed structure is as shown in Figure 19
-D.5 Integration Prompt
+F.5 Integration Prompt
 This stage integrates the outputs from previous modules into a unified final response. The detailed prompt is as shown in Figure 20
-D.6 BoN Ranking Prompt
-In this step, we employ a Best-of-N (BoN) ranking mechanism to select the most appropriate final output from multiple candidates generated by the model.
-The ranking prompt used to guide this selection process is shown in Figure 18.
+Figure 10: Reasoning pattern comparison on an AIME 2024 problem (ground truth: 294).
+Blue highlights indicate where SafR and Qwen3-8B share identical reasoning structures.
+Red highlights show Star-1’s divergent approach leading to an incorrect answer (588).
 Figure 11: Example of Benign Reframing.
@@
 Figure 20: Prompt structure used in the Safety Backtrack stage, incorporating contextual transition phrases based on classification results.
-Figure 18: Prompt structure used in the Best-of-N (BoN) ranking stage to select the best output from multiple candidates.